### v1

# EEG-SCMM: Soft Contrastive Masked Modeling for Cross-Corpus EEG-Based Emotion Recognition

Abstract
Emotion recognition using electroencephalography (EEG) signals has garnered widespread attention in recent years. However, existing studies have struggled to develop a sufficiently generalized model suitable for different datasets without re-training (cross-corpus). This difficulty arises because distribution differences across datasets far exceed the intra-dataset variability. To solve this problem, we propose a novel Soft Contrastive Masked Modeling (SCMM) framework. Inspired by emotional continuity, SCMM integrates soft contrastive learning with a new hybrid masking strategy to effectively mine the ”short-term continuity” characteristics inherent in human emotions. During the self-supervised learning process, soft weights are assigned to sample pairs, enabling adaptive learning of similarity relationships across samples. Furthermore, we introduce an aggregator that weightedly aggregates complementary information from multiple close samples based on pairwise similarities among samples to enhance fine-grained feature representation, which is then used for original sample reconstruction. Extensive experiments on the SEED, SEED-IV and DEAP datasets show that SCMM achieves state-of-the-art (SOTA) performance, outperforming the second-best method by an average accuracy of 4.26% under two types of cross-corpus conditions (same-class and different-class) for EEG-based emotion recognition.
Introduction
Emotions are human attitudinal experiences and behavioral responses to objective things, closely related to health conditions and behavioral patterns (Wang, Zhang, and Tang 2024). Compared to speech (Singh and Goel 2022), gestures (Noroozi et al. 2018), and facial expressions (Canal et al. 2022), electroencephalography (EEG) offers a more direct and objective measurement of human emotions by capturing brain activity across various scalp locations (Hu et al. 2019). Recently, researchers have increasingly emphasised EEG-based emotion recognition (Zhong, Wang, and Miao 2020; Zhao, Yan, and Lu 2021; Zhang, Liu, and Zhong 2022; Gao et al. 2024), aiming to advance the development of affective brain-computer interfaces (aBCIs). However, three critical challenges remain to be addressed in current approaches.
(1) Dataset Specificity. Most existing EEG-based emotion recognition methods are typically designed for a single dataset, necessitating model retraining when the dataset changes. This requirement significantly limits the model’s generalizability and scalability, hindering its application across different datasets. To tackle this issue, the concept of cross-corpus has been proposed, which is designed to be generalized across multiple datasets. A cross-corpus model is trained on one dataset and can be directly applied to another without the need for retraining from scratch. This concept, which originated in natural language processing (Schuller et al. 2010; Zhang et al. 2011), has been extended to various domains in recent years (Rayatdoost and Soleymani 2018; Chien, Yang, and Lee 2020; Ryumina, Dresvyanskiy, and Karpov 2022). Although existing EEG-based emotion recognition methods, such as BiDANN (Li et al. 2018), TANN (Li et al. 2021), and PR-PL (Zhou et al. 2023), have achieved superior performance in within-subject or cross-subject tasks, their efficacy significantly degrades in cross-corpus scenarios, where differences in data distribution across datasets far exceed the variability within a single dataset (Rayatdoost and Soleymani 2018).
Figure 1: (a) An illustration of emotional continuity. We take the sample $x_{i}$ at second $t$ within an EEG trial as the anchor, and calculate the cosine similarity between $x_{i}$ and its neighboring sample $x_{j}$, as well as the distant sample $x_{k}$. High cosine similarities indicate that human emotions remain relatively stable and similar over a certain period. (b) Hard CL vs. SCMM. Traditional hard contrastive learning considers the embeddings $z_{i}$ and $\widetilde{z}_{i}$ of the same sample $x_{i}$ and its augmented view $\widetilde{x}_{i}$ as positive pairs, while different samples and their augmented views are treated as negative pairs. When computing the contrastive loss, the weights for positives and negatives are set to 1 and 0, respectively. In contrast, the soft contrastive learning strategy designed in SCMM generates soft assignments for different sample pairs based on their distances in the original data space.
(2) Data Availability. Current approaches for cross-domain or cross-corpus EEG-based emotion recognition rely heavily on domain adaptation techniques, which depend extensively on the availability of labeled source data and unlabeled target data. For example, AD-TCN (He, Zhong, and Pan 2022) learned an asymmetric mapping that adapts the target domain feature encoder to the source domain, aiming to eliminate the complicated steps of target domain labeling and improve the model performance in cross-domain scenarios. Similarly, E2STN (Zhou et al. 2024) integrated content information from the source domain with style information from the target domain to create stylized emotional EEG representations. However, these methods require prior access to all labeled source data and unlabeled target data for model training, presenting a significant limitation due to data availability constraints (Liu et al. 2024a).
(3) Ignorance of Emotional Continuity. Unlike domain adaptation techniques, contrastive learning (CL) achieves superior performance without relying on labeled data, and has demonstrated significant potential in enhancing feature representation capabilities across various domains (Chen et al. 2020; Radford et al. 2021; Zhang et al. 2022). Current CL-based methods for EEG-based emotion recognition consider an anchor and its augmented views as positive pairs, while treating all other samples as negatives, as shown in Fig. 1(b) (Hard CL). For example, CLISA (Shen et al. 2022) employed contrastive learning to minimize the inter-subject differences by maximizing the similarity in EEG representations across subjects. JCFA (Liu et al. 2024a) performed joint contrastive learning across three domains to align the time- and frequency-based embeddings of the same sample in the latent time-frequency space, achieving state-of-the-art (SOTA) performance in cross-corpus scenarios. However, psychological and neuroscientific studies have shown that emotions exhibit significant ”short-term continuity” characteristics (Davidson 1998; Houben, Van Den Noortgate, and Kuppens 2015). In other words, human emotions are relatively stable over certain periods, with sudden changes being rare. As illustrated in Fig. 1(a), a high cosine similarity is maintained between an anchor sample $x_{i}$ and its neighboring sample $x_{j}$, and even a distant sample $x_{k}$ separated by extended periods (e.g., 60 seconds). Given this nature of emotions, we propose that the definition of positive pairs in CL-based EEG analysis should extend beyond just the anchor and its augmented views. Instead, it should include a broader range of similar samples, especially those that are temporally proximal, as shown in Fig. 1(b) (SCMM). However, existing methods like JCFA (Liu et al. 2024a), which follow the traditional CL paradigm (Chen et al. 2020), may incorrectly pull apart similar but not identical samples, thus failing to capture the inherent correlations of EEG signals.
To address the aforementioned three critical issues, we propose a novel Soft Contrastive Masked Modeling (SCMM) framework for cross-corpus EEG-based emotion recognition. Unlike traditional hard CL shown in Fig. 1(b), SCMM considers emotional continuity and incorporates soft assignments of sample pairs. This approach enables the model to identify the fine-grained relationships between samples in a self-supervised manner, thereby enhancing the generalizability of EEG representations. Comprehensive experiments on three well-recognized datasets demonstrate that SCMM consistently achieves SOTA performance, highlighting its superior capability and stability. In summary, the main contributions of SCMM are outlined as follows:
  • Inspired by the nature of emotions, we propose a novel SCMM framework to address cross-corpus EEG-based emotion recognition. This approach assigns soft weights to sample pairs during contrastive learning to capture the similarity relationships between different samples. As a result, better feature representations of EEG signals are learned in a self-supervised manner. To the best of our knowledge, this is the first study to introduce soft contrastive learning into EEG-based emotion recognition.
  • We also develop a new hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships, which is essential for enhancing contrastive learning. Additionally, we introduce an aggregator that weightedly aggregates complementary information from the embeddings of multiple close samples, enabling fine-grained feature learning and improving the model’s overall capability.
  • We conduct extensive experiments on three well-known datasets (SEED, SEED-IV, and DEAP), showing that SCMM achieves consistent SOTA performance compared to eight competitive methods, surpassing the second-best method by an average accuracy of 4.26%.
Methodology
Figure 2: The overall framework of SCMM. The pre-training process of SCMM involves three modules: (1) hybrid masking, (2) soft contrastive learning, and (3) aggregate reconstruction.
Problem Definition
Given an unlabeled pre-training EEG emotion dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$ with $N$ samples, where each sample $x_{i}\in\mathbb{R}^{C\times F}$ contains $C$ channels and $F$-dimensional features, the goal is to learn a nonlinear embedding function $f_{\theta}$. This function is designed to map $x_{i}$ to its representation $h_{i}$ that best describes itself by leveraging the emotional continuity inherent in EEG signals. Ultimately, the pre-trained model is capable of producing generalizable EEG representations that can be effectively used across different EEG emotion datasets.
Model Architecture
The overall framework of SCMM is illustrated in Fig. 2, which includes three main modules: hybrid masking, soft contrastive learning, and aggregate reconstruction. Below, we detail the specific design of each module and the pre-training process of SCMM.
Hybrid Masking
The selection of masking strategies is crucial for CL and masked modeling (Zhang et al. 2024; Liu et al. 2024b). For an input EEG sample $x_{i}\in\mathcal{X}$, most existing methods use random masking (Zhang, Liu, and Zhong 2022) or channel masking (Li et al. 2022) to generate the masked sample $\widetilde{x}_{i}$. The random masking strategy masks samples along the feature dimension, ignoring the inter-channel relationships of EEG signals. While a large masking ratio (e.g., 75%) can mask entire portions of certain channels, it complicates the modeling process due to significant information loss. Conversely, the channel masking strategy masks features across all dimensions of the selected channels, losing the relationships between different dimensional features. Neither approach captures both channel and feature relationships simultaneously. Therefore, we develop a new hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships.
Specifically, we first generate a random masking matrix ${\rm Mask}_{R}\in\{0,1\}$ with dimensions $C\times F$, and a channel masking matrix ${\rm Mask}_{C}\in\{0,1\}$ with dimensions $C\times F$, both derived from binomial distributions with masking ratios $r\in(0,1)$. Here, the element values in each row of ${\rm Mask}_{C}$ are either all 1s or all 0s. Next, we generate a probability matrix $U\in[0,1]$ with dimensions $C\times 1$ for hybrid masking, which is drawn from a uniform distribution. The hybrid masking process is defined as:
$\widetilde{x}_{i,c}=\begin{cases}x_{i,c}\odot{\rm Mask}_{R,c}&\text{if }\mu<U_{c}\leq 1\\ x_{i,c}\odot{\rm Mask}_{C,c}&\text{if }0\leq U_{c}\leq\mu\end{cases},$ (1)
where $x_{i,c}$ represents the $c$-th channel of $x_{i}$, and $\widetilde{x}_{i,c}$ is the corresponding masked sample. $\odot$ denotes element-wise multiplication. $U_{c}$ is the probability value in the $c$-th row, and $\mu$ is a probability threshold that controls the weights of the two masking strategies. By integrating the hybrid masking strategy in SCMM, we enhance the diversity of masked samples, encouraging the model to learn richer and more robust feature representations that account for both channel and feature relationships within EEG signals. Figure 3 illustrates the differences between three masking strategies.
Figure 3: Illustrations of different masking strategies. (a) compares various masking strategies, and (b) presents examples of generated masked samples using three strategies. The masking ratio and threshold are set to $r=0.5$ and $\mu=0.1$, respectively.
Soft Contrastive Learning
Traditional hard CL treats augmented views generated from the same sample as positive pairs, and those from different samples as negative pairs (Chen et al. 2020; Eldele et al. 2021; Yue et al. 2022). During the computation of the contrastive loss, hard values (1 or 0) are assigned to sample pairs, as illustrated in Fig. 1(b) (Hard CL). However, we argue that this approach fails to account for the “short-term continuity” characteristic inherent in human emotions, leading to inaccurate modeling of inter-sample relationships and hindering the generalizability of the learned embeddings.
To address this issue, we propose defining soft assignments for different sample pairs, as shown in Fig. 1(b) (SCMM). We first input $x_{i}$ and $\widetilde{x}_{i}$ into an encoder $E$ that maps samples to embeddings, denoted as $h_{i}=E(x_{i})$ and $\widetilde{h}_{i}=E(\widetilde{x}_{i})$. These embeddings are then projected into a latent space $\mathcal{Z}$ using a projector $P$, resulting in $z_{i}=P(h_{i})$ and $\widetilde{z}_{i}=P(\widetilde{h}_{i})$. Next, we perform soft contrastive learning in $\mathcal{Z}$ using $z_{i}$ and $\widetilde{z}_{i}$. Specifically, for a given pair of samples ($x_{i}$, $x_{j}$), we first calculate the normalized distance $D(x_{i},x_{j})$ between $x_{i}$ and $x_{j}$ in the original data space as:
$D(x_{i},x_{j})=Norm(Dist(x_{i},x_{j}))\in[0,1],$ (2)
where $Dist(\cdot,\cdot)$ is a metric function used to measure the distance between sample pairs, and $Norm(\cdot)$ denotes min-max normalization. In the experiments, we take the negative of cosine similarity as the metric function. Based on the normalized distance $D(x_{i},x_{j})$, we then define a soft assignment $w{(x_{i},x_{j})}$ for each pair of samples ($x_{i}$, $x_{j}$) using the sigmoid function $\sigma(x)=1/(1+{\rm exp}(-x))$:
$w{(x_{i},x_{j})}=2\alpha\cdot\sigma(-D(x_{i},x_{j})/\tau_{w}),$ (3)
where $\alpha\in[0,1]$ is a boundary parameter that controls the upper bound of soft assignments. $\tau_{w}$ is a sharpness parameter, where smaller values of $\tau_{w}$ result in greater differences in $w(\cdot,\cdot)$ between sample pairs, and vice versa. Figure 4 depicts heat maps of soft assignments $w(\cdot,\cdot)$ under different sharpness parameters $\tau_{w}$.
Figure 4: Soft assignments $w(\cdot,\cdot)$ with different $\tau_{w}$. We use heat maps to visualize the soft assignments with different sharpness parameters $\tau_{w}$. For clarity, we set the upper bound $\alpha=1$ in the figure. Best viewed in color.
Leveraging the soft assignments for all sample pairs, we propose a soft contrastive loss to refine the traditional hard contrastive loss. Specifically, for a pair of projected embeddings $(z_{i},\widetilde{z}_{i})$, we first calculate the softmax probability of the relative similarity among all similarities as:
$p(z_{i},\widetilde{z}_{i})=\frac{{\rm exp}({\rm sim}(z_{i},\widetilde{z}_{i})/\tau_{c})}{\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{c})},$ (4)
where ${\rm sim(\cdot,\cdot)}$ refers to the cosine similarity, and $\tau_{c}$ is a temperature parameter used to adjust the scale. Based on $p(z_{i},\widetilde{z}_{i})$, the soft contrastive loss is then defined as:
$\begin{split}\mathcal{L}_{C,i}&=-{\rm log}\thinspace p(z_{i},\widetilde{z}_{i})\\ &-\sum_{\begin{subarray}{c}x^{\prime}\in\Omega\backslash\{x_{i},\widetilde{x}_{i}\}\\ z^{\prime}\in\mathcal{Z}\backslash\{z_{i},\widetilde{z}_{i}\}\end{subarray}}w{(x_{i},x^{\prime})}\cdot{\rm log}\thinspace p(z_{i},z^{\prime}),\end{split}$ (5)
where $\Omega=\mathcal{X}\cup\widetilde{\mathcal{X}}$ represents the union of the data spaces of the original and masked samples. By assigning soft weights to different sample pairs, the model is encouraged to better capture the inherent correlations across different samples. During the training process, the final soft contrastive loss $\mathcal{L}_{C}$ is computed by summing and averaging $\mathcal{L}_{C,i}$ across all samples within a mini-batch. Notably, when $\forall{w(x_{i},x^{\prime})}=0$, the soft contrastive loss reduces to the traditional hard contrastive loss.
Aggregate Reconstruction
To further capture the fine-grained relationships between different samples, we incorporate an aggregator for weighted aggregation and reconstruction. Current approaches for masked EEG modeling typically reconstruct the masked portion based on the unmasked portion of a single masked sample (Lan et al. 2024; Pang et al. 2024), following the learning paradigm of MAE (He et al. 2022). However, this single-sample reconstruction strategy overlooks the interactions between samples, leading to a complex and inadequate reconstruction process.
To overcome this limitation, we introduce an aggregator that improves the traditional single-sample reconstruction process. Specifically, we first calculate the cosine similarity between each pair of projected embeddings $(z_{i},z_{j})$ within a mini-batch, resulting in a similarity matrix $\mathcal{S}$. Based on the pairwise similarities in $\mathcal{S}$, the aggregator then performs weighted aggregation of the embedding $h_{i}$. The weighted aggregation process is defined as:
$h_{i}^{r}=\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}(\frac{{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{c})}{\sum_{z^{\prime\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime\prime})/\tau_{c})}\cdot h^{\prime}),$ (6)
where $h^{\prime}\in\mathcal{H}\backslash\{h_{i}\}$ represents the encoded embedding corresponding to the projected embedding $z^{\prime}$, and $\mathcal{H}$ denotes the embedding space of the encoder $E$. This approach allows for a more comprehensive reconstruction by aggregating complementary information and incorporating similar features from different samples during the reconstruction process, while suppressing interference from irrelevant noise samples. Finally, the reconstructed embedding $h_{i}^{r}$ is fed into a lightweight decoder $D$ to obtain the reconstructed sample $x_{i}^{r}$. Following the masked modeling paradigm, we use Mean Square Error (MSE) as the reconstruction loss for model optimization, which is defined as:
$\mathcal{L}_{R,i}=\|x_{i}-x_{i}^{r}\|_{2}^{2}.$ (7)
Similar to the soft contrastive loss $\mathcal{L}_{C}$, the final reconstruction loss $\mathcal{L}_{R}$ is computed by summing and averaging $\mathcal{L}_{R,i}$ across all samples within a mini-batch.
The pre-training process of SCMM
During the pre-training process, SCMM is trained by jointly optimizing $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$. The overall pre-training loss is defined as:
$\mathcal{L}=\lambda_{C}\mathcal{L}_{C}+\lambda_{R}\mathcal{L}_{R},$ (8)
where $\lambda_{C}$ and $\lambda_{R}$ are trade-off hyperparameters, which are adaptively adjusted according to the homoscedastic uncertainty of each loss item. Algorithm 1 details the pre-training process of the proposed SCMM.
Algorithm 1 The pre-training process of SCMM.
Input: Unlabeled pre-training dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$. The number of pre-training epochs.
The pre-training process:
1:  Randomly initialize the model parameters $\theta$;
2:  for $epoch=1$ to epochs do
2:      /*All operations are performed within a mini-batch*/
3:     Generate the masked sample $\widetilde{x}_{i}$ of each input EEG sample $x_{i}$ using hybrid masking in Eq. (1);
4:     Generate $h_{i}$ and $\widetilde{h}_{i}$ by feeding $x_{i}$ and $\widetilde{x}_{i}$ into $E$;
5:     Generate $z_{i}$ and $\widetilde{z}_{i}$ by feeding $h_{i}$ and $\widetilde{h}_{i}$ into $P$;
6:     Compute the normalized distance $D(x_{i},x_{j})$ for each pair of samples $(x_{i},x_{j})$ using Eq. (2);
7:     Generate the soft assignment $w(x_{i},x_{j})$ for each pair of samples $(x_{i},x_{j})$ using Eq. (3);
8:     Compute the soft contrastive loss $\mathcal{L}_{C}$ using Eq. (5);
9:     Compute the pairwise cosine similarity for each pair of projected embeddings $(z_{i},z_{j})$;
10:     Generate the reconstructed embedding $h_{i}^{r}$ of each $h_{i}$ through weighted aggregation in Eq. (6);
11:     Reconstruct $x_{i}^{r}$ by feeding $h_{i}^{r}$ into $D$;
12:     Compute the reconstruction loss $\mathcal{L}_{R}$ using Eq. (7);
13:     Compute the pre-training loss $\mathcal{L}$ using Eq. (8);
14:     Update the model parameters $\theta$;
15:  end for
16:  return  The pre-trained SCMM model $f_{\theta}$.
Experiments
Methods Same-Class Different-Class
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
BiDANN (Li et al. 2018) 49.24 / 10.49 60.46 / 11.17 - -
TANN (Li et al. 2021) 58.41 / 07.16 60.75 / 10.61 - -
PR-PL (Zhou et al. 2023)* 61.01 / 10.55 58.74 / 10.71 - -
E2STN (Zhou et al. 2024) 60.51 / 05.41 61.24 / 15.14 - -
SimCLR (Tang et al. 2021)* 47.27 / 08.44 46.89 / 13.41 44.19 / 09.28 42.03 / 10.05
Mixup (Wickstrøm et al. 2022)* 56.86 / 16.83 55.70 / 16.28 54.55 / 17.95 45.79 / 15.16
MAE (He et al. 2022)* 86.49 / 10.57 83.87 / 08.53 86.02 / 08.96 76.74 / 09.18
JCFA (Liu et al. 2024a)* 67.53 / 12.36 62.40 / 07.54 65.99 / 14.04 52.67 / 05.86
SCMM (Ours) 91.61 / 07.56 (+05.12) 87.24 / 08.35 (+03.37) 91.26 / 07.91 (+05.24) 80.89 / 08.69 (+04.15)
Table 1: Experimental results on SEED and SEED-IV under two cross-corpus conditions: same-class and different-class. ”*” indicates that the results are reproduced by ourselves. A $\rightarrow$ B denotes that A is the pre-training dataset, while B is the dataset for model fine-tuning and testing. Best results are highlighted in bold, while the second-best results are underlined.
Hard $\mathcal{L}_{C}$ Soft $\mathcal{L}_{C}$ $\mathcal{L}_{R}$ Same-Class Different-Class
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
90.08 / 09.24 84.43 / 11.83 90.26 / 08.71 77.79 / 08.04
90.73 / 08.48 85.07 / 11.05 90.96 / 08.36 78.32 / 07.19
89.68 / 09.32 84.24 / 11.90 89.45 / 09.10 77.24 / 09.14
90.30 / 07.94 85.95 / 08.74 90.91 / 08.61 79.82 / 07.00
91.61 / 07.56 87.24 / 08.35 91.26 / 07.91 80.89 / 08.69
Table 2: Ablation study on SEED and SEED-IV under two cross-corpus conditions: same-class and different-class.
Datasets
We conduct extensive experiments on three publicly available datasets, SEED (Zheng and Lu 2015), SEED-IV (Zheng et al. 2018), and DEAP (Koelstra et al. 2011), to evaluate the model performance of SCMM. These datasets are diverse in terms of EEG equipment, emotional stimuli, data specifications, labeling approaches and subjects, making them well-suited for assessing the model’s efficacy in cross-corpus EEG-based emotion recognition tasks. In our experiments, we use differential entropy (DE) features as inputs. Detailed descriptions of the datasets and pre-processing procedures are provided in Appendix A.
Implementation Details
In the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking. We set $\alpha$ to 0.5, $\tau_{w}$ to 0.05, and $\tau_{c}$ to 0.5 for soft CL. The Adam optimizer is utilized with a learning rate of 5e-4 and a weight decay of 3e-4. The pre-training process is conducted over 200 epochs with a batch size of 256. We save the parameters $\theta$ from the final epoch as the pre-trained model. In the fine-tuning stage, we input the embeddings $h_{i}$ generated by the encoder $E$ into an emotion classifier consisting of a 2-layer MLP for final emotion recognition. The classifier is optimized using cross-entropy (CE) loss, and the fine-tuning process is conducted over 50 epochs with a batch size of 128. All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce RTX 3090 GPU. Further implementation details and hyperparameter analysis can be found in Appendix B and Appendix E.
Baselines and Experimental Setup
We compare the proposed SCMM against eight competitive baseline methods, including four conventional deep learning methods: BiDANN (Li et al. 2018), TANN (Li et al. 2021), PR-PL (Zhou et al. 2023), and E2STN (Zhou et al. 2024), as well as four self-supervised learning models: SimCLR (Chen et al. 2020; Tang et al. 2021), Mixup (Zhang et al. 2018; Wickstrøm et al. 2022), MAE (He et al. 2022), and JCFA (Liu et al. 2024a). Notably, E2STN and JCFA are specifically designed for cross-corpus EEG-based emotion recognition. To ensure a fair comparison, we adopt a cross-corpus subject-independent protocol in our experiments, consistent with the setup used by JCFA. We evaluate the model performance using the average accuracy and standard deviation (ACC / STD %) across all subjects in the test set. More details about baseline methods and experimental settings are provided in Appendix C.
Experimental Results
Methods DEAP $\rightarrow$ SEED3 SEED3 $\rightarrow$ DEAP (Valence) SEED3 $\rightarrow$ DEAP (Arousal)
SimCLR (Tang et al. 2021)* 53.12 / 13.12 53.75 / 03.61 51.79 / 04.54
Mixup (Wickstrøm et al. 2022)* 48.75 / 14.37 60.62 / 08.68 60.11 / 07.69
MAE (He et al. 2022)* 83.69 / 10.10 72.19 / 07.24 70.50 / 06.30
JCFA (Liu et al. 2024a)* 64.69 / 12.28 61.59 / 06.26 61.06 / 07.37
SCMM (Ours) 91.70 / 08.07 (+08.01) 73.96 / 06.75 (+01.77) 72.66 / 05.67 (+02.16)
Table 3: Experimental results for cross-corpus EEG-based emotion recognition on SEED and DEAP.
Evaluation on the SEED and SEED-IV datasets
To fully validate the model performance of SCMM, we conduct extensive experiments under two cross-corpus conditions: (1) same-class and (2) different-class.
(1) For the same-class cross-corpus validation, we conduct two experiments on the SEED and SEED-IV 3-category datasets: pre-training on SEED-IV and fine-tuning on SEED (SEED-IV3 $\rightarrow$ SEED3), and pre-training on SEED and fine-tuning on SEED-IV (SEED3 $\rightarrow$ SEED-IV3). In both experiments, all samples corresponding to fear emotions in the SEED-IV dataset are excluded. The left two columns in Table 1 present the comparison results, indicating that SCMM achieves SOTA performance in both experiments. Specifically, our model achieves classification accuracies of 91.61% and 87.24% with standard deviations of 7.56% and 8.35% in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments, outperforming the second-best method MAE by accuracies of 5.12% and 3.37%, respectively.
(2) For the different-class cross-corpus validation, we conduct another two experiments, denoted as SEED-IV4 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV4. These experiments aim to evaluate the model performance when the pre-training and fine-tuning datasets contain different emotion categories. Experimental results in the right two columns of Table 1 demonstrate that SCMM achieves the best performance in both experiments. Specifically, our model achieves classification accuracies of 91.26% and 80.89% with standard deviations of 7.91% and 8.69% in the SEED-IV4 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV4 experiments, surpassing the second-best method MAE by 5.24% and 4.15% in accuracies, respectively. Note that the traditional CL-based models SimCLR, Mixup, and JCFA exhibit relatively poor performance, primarily due to their use of raw EEG signals as inputs. In addition, SCMM significantly outperforms the traditional masked modeling framework MAE, highlighting the superiority of SCMM. In summary, extensive experimental results on the SEED and SEED-IV datasets confirm that our model exhibits superior performance in multiple cross-corpus EEG-based emotion recognition tasks.
Ablation study
To assess the validity of each module in SCMM, we conduct a comprehensive ablation study on the SEED and SEED-IV datasets. Specifically, we design five different models below. (1) Hard $\mathcal{L}_{C}$ trains the model with only hard contrastive loss. (2) Soft $\mathcal{L}_{C}$ trains the model with only soft contrastive loss. (3) $\mathcal{L}_{R}$ trains the model with only reconstruction loss. (4) Hard $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$ trains the model using both hard contrastive loss and reconstruction loss. (5) Soft $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$ trains the model using both soft contrastive loss and reconstruction loss, representing the full model.
Table 2 presents the results of ablation experiments. Specifically, (1) Hard $\mathcal{L}_{C}$ and (2) Soft $\mathcal{L}_{C}$: the first two models compare the performance of hard CL and soft CL. Experimental results indicate that our well-designed soft contrastive learning effectively improves the classification accuracy. (3) $\mathcal{L}_{R}$: the third model removes the soft contrastive loss and trains the model with only reconstruction loss. The results show that the model performs worst in the absence of the contrastive learning constraint. (4) Hard $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$ and (5) Soft $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$: the last two models compare the performance when combining hard and soft contrastive loss with reconstruction loss, respectively. Experimental results demonstrate that the full model achieves the best performance in all experiments, indicating that SCMM significantly enhances the model performance and stability by combining soft contrastive learning and aggregate reconstruction. This improvement is evident under different cross-corpus conditions, demonstrating the feasibility of extending SCMM to real-life aBCI applications.
Discussion
Method ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
OS 91.61 / 07.56 87.24 / 08.35 91.25 / 07.91 80.89 / 08.69
ES 89.99 / 10.25 85.75 / 14.00 90.31 / 08.59 79.04 / 06.95
DEAP $\rightarrow$ SEED3 SEED3 $\rightarrow$ DEAP (Valence) SEED3 $\rightarrow$ DEAP (Arousal) -
OS 91.70 / 08.01 73.96 / 06.75 72.66 / 05.67 -
ES 90.64 / 07.97 72.75 / 07.06 71.58 / 05.72 -
Table 4: Comparison of soft contrastive learning in the original data space (OS) and embedding space (ES).
Generalization capability analysis
To further validate the generalization capability of SCMM, we conduct additional experiments on the SEED and DEAP datasets under a different-class cross-corpus scenario, denoted as DEAP $\rightarrow$ SEED3, SEED3 $\rightarrow$ DEAP (Valence) and SEED3 $\rightarrow$ DEAP (Arousal). It is notable that the EEG acquisition equipment, emotional stimuli, data specifications, labeling approaches and subjects are completely different between the two datasets. Table 3 presents the experimental results of SCMM compared to existing methods. Specifically, for the DEAP $\rightarrow$ SEED3 experiment, SCMM achieves an accuracy of 91.70% with a standard deviation of 8.07%, outperforming the second-best method MAE by an accuracy of 8.01%. For the SEED3 $\rightarrow$ DEAP (Valence) and SEED3 $\rightarrow$ DEAP (Arousal) experiments, our model achieves classification accuracies of 73.96% and 72.66% with standard deviations of 6.75% and 5.67%, surpassing the second-best method MAE by 1.77% and 2.16% in accuracies. Experimental results demonstrate that the proposed SCMM maintains excellent performance even when the pre-training and fine-tuning datasets are completely different, highlighting its superior generalization capability.
Model performance with limited fine-tuning data
Figure 5: Model performance with limited labeled data for fine-tuning on SEED and SEED-IV.
We further explore the model performance of SCMM on the SEED and SEED-IV datasets when fine-tuning labeled data is limited. Specifically, we randomly select 1%, 5%, 10% and 20% of labeled samples from the fine-tuning dataset for model fine-tuning, while the remaining samples are used for testing. Figure 5 shows the classification accuracy curves. For the SEED-IV3 $\rightarrow$ SEED3 and SEED-IV4 $\rightarrow$ SEED3 experiments, SCMM achieves classification accuracies exceeding 70% with only 1% of labeled data. The accuracies significantly improve as the proportion of labeled samples increases, reaching close to 100% with 10% of labeled data. Meanwhile, our model achieves classification accuracies over 60% and 70% with only 1% of labeled data in the SEED3 $\rightarrow$ SEED-IV3 and SEED3 $\rightarrow$ SEED-IV4 experiments. The classification accuracies exceed 90% when fine-tuning with 10% of labeled samples in both experiments. In summary, experimental results indicate that SCMM maintains superior performance even with limited labeled data for fine-tuning, showing its outstanding robustness and potential in few-shot scenarios. The complete results and computational complexity analysis are provided in Appendix F.
Comparison of different masking strategies
Strategies SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
Random 90.30 / 08.80 84.63 / 10.99
Channel 90.25 / 08.68 85.91 / 10.97
Hybrid 91.61 / 07.56 87.24 / 08.35
Table 5: Comparative experiments of different masking strategies on the SEED and SEED-IV 3-category datasets.
To assess the impact of various masking strategies, we conduct comparative experiments on the SEED and SEED-IV 3-category datasets using three strategies: random, channel, and hybrid masking. Table 5 presents the experimental results, showing that the hybrid masking strategy achieves the highest accuracy and the lowest standard deviation in both experiments. This suggests that the integration of different masking approaches significantly improves the model performance and stability. More details on the different masking strategies are presented in Appendix D.
Comparison of soft CL in the original data space and embedding space
While soft contrastive learning has been explored across various domains, most methods focus on computing soft assignments for contrastive loss in the embedding space (Dwibedi et al. 2021; Yèche et al. 2021). However, we argue that utilizing similarities in the original data space offers superior self-supervision and is particularly well-suited for EEG emotional data. To validate this, we conduct additional experiments on the SEED, SEED-IV and DEAP datasets to verify the effectiveness of soft contrastive learning in the original data space. Specifically, we modify the metric function $Dist(\cdot,\cdot)$ to use similarities between projected embeddings, shifting the computation of soft assignments from the original data space to the embedding space. Table 4 presents the experimental results, demonstrating that soft contrastive learning in the original data space consistently outperforms the embedding space in all experiments. Furthermore, this approach allows for pre-computing cosine similarities between original sample pairs offline, thus reducing computational resource requirements and improving training efficiency.
Visualization
Figure 6: (a) Heat maps of pairwise similarity matrices. (b) t-SNE visualization of the learned embeddings. Best viewed in color. Zoom in for better view.
To evaluate whether sample-wise relationships are preserved in the encoder, we randomly select 100 test samples from the SEED dataset, and visualize the pairwise cosine similarity between sample pairs. Additionally, we select all test samples of one subject from the SEED dataset and visualize the learned embeddings of SCMM using t-SNE (Van der Maaten and Hinton 2008). Figure 6(a) presents heat maps of pairwise similarity matrices, where darker colors indicate higher similarity between samples. Traditional hard CL identifies only coarse-grained relationships across samples from different emotion categories, especially for the most challenging-to-recognize negative and neutral emotions. In contrast, SCMM effectively captures the fine-grained relationships between samples of different categories. Moreover, the results of t-SNE visualization in Fig. 6(b) indicate that the proposed SCMM better clusters samples within the same category and increases the inter-class distance compared to traditional hard CL, thus enhancing the classification performance. More visualization results can be found in Appendix G.
Conclusion
This paper proposes a novel self-supervised pre-training framework, Soft Contrastive Masked Modeling (SCMM), for cross-corpus EEG-based emotion recognition. Unlike traditional contrastive learning models, SCMM integrates soft contrastive learning with a hybrid masking strategy to effectively capture the ”short-term continuity” characteristics inherent in human emotions, and produce stable and generalizable EEG representations. Additionally, an aggregator is developed to weightedly aggregate complementary information from multiple close samples, thereby enhancing fine-grained feature representation capability in the modeling process. Extensive experiments on three well-recognized datasets show that SCMM consistently achieves SOTA performance in cross-corpus EEG-based emotion recognition tasks under both same-class and different-class conditions. Comprehensive ablation study and parameter analysis confirm the superior performance and robustness of SCMM. Visualization results indicate that our model effectively reduces the distance between similar samples within the same category, and captures more fine-grained relationships across samples. These findings suggest that SCMM enhances the feasibility of extending the proposed method to real-life aBCI applications.
Appendix A Datasets
A.1 Dataset Description
We conduct extensive experiments on three well-recognized datasets, SEED (Zheng and Lu 2015), SEED-IV (Zheng et al. 2018), and DEAP (Koelstra et al. 2011), to evaluate the model performance of SCMM in cross-corpus EEG-based emotion recognition tasks. Table 6 provides a detailed description of the three datasets.
Datasets Subjects Sessions $\times$ Trials Channels Sampling Rate Classes
SEED 15 3 $\times$ 15 62 1kHz 3 (Negative, Neutral, Positive)
SEED-IV 15 3 $\times$ 15 62 1kHz 4 (Sad, Neutral, Fear, Happy)
DEAP 32 1 $\times$ 40 32 512Hz Valence: 1 - 9, Arousal: 1 - 9
Table 6: Detailed description of the experimental datasets.
(1) SEED (Zheng and Lu 2015) was developed by Shanghai Jiao Tong University. The dataset used a 62-channel ESI NeuroScan System based on the international 10-20 system to record EEG signals from 15 subjects (7 males and 8 females) under different video stimuli at a sampling rate of 1kHz. Each subject participated in 3 sessions. In each session, each subject was required to watch 15 movie clips consisting of 3 different emotional states: negative, neutral and positive. Each emotional state contains a total of 5 movie clips, corresponding to 5 trials.
(2) SEED-IV (Zheng et al. 2018) used the same EEG acquisition equipment as the SEED dataset, but with different video stimuli, emotion categories and subjects. The dataset recorded EEG signals from 15 subjects under different video stimuli at a sampling rate of 1kHz. Each subject participated in 3 sessions. In each session, each subject was required to watch 24 movie clips containing 4 different emotions: sad, neutral, fear and happy. Each emotion contains a total of 6 movie clips, corresponding to 6 trials.
(3) DEAP (Koelstra et al. 2011) was constructed by Queen Mary University of London. The dataset used a 128-channel Biosemi ActiveTwo System to record EEG signals from specific 32 channels of 32 subjects (16 males and 16 females) while watching 40 one-minute music videos at a sampling rate of 512Hz. The 40 one-minute videos elicited different emotions according to the valence-arousal dimension. Specifically, the valence-arousal emotional model, first proposed by Russell (Russell 1980), places each emotional state on a two-dimensional scale. The first dimension represents valence, ranging from negative to positive, and the second dimension represents arousal, ranging from calm to exciting. Participants rated valence and arousal using a continuous scale of 1 to 9 after watching each video clip.
A.2 Pre-processing Procedures
For the SEED and SEED-IV datasets, the raw EEG signals were initially downsampled to 200Hz and filtered through a bandpass filter of 0.3-50Hz to filter noise and remove artifacts. Then, the data were divided into multiple non-overlapping segments using sliding windows of 1 second for SEED and 4 seconds for SEED-IV, respectively. After that, we extracted differential entropy (DE) features for each channel of each segment at five frequency bands: Delta (1-4Hz), Theta (4-8Hz), Alpha (8-14Hz), Beta (14-31Hz), and Gamma (31-50Hz). Finally, the DE features from 62 channels and 5 bands were formed into a feature matrix of shape $62\times 5$, which serves as input to the SCMM model. The extraction of DE features can be expressed as:
$DE(x)=\frac{1}{2}{\rm log}(2\pi e\sigma^{2}),$ (9)
Here, $x$ represents an EEG signal segment of a specific length that approximately obeys a Gaussian distribution $N(\mu,\sigma^{2})$, where $\sigma$ denotes the standard deviation of $x$, and $e$ is the Euler constant.
For the DEAP dataset, the raw EEG signals were first downsampled to 128Hz and denoised by a bandpass filter of 4-45Hz. Subsequently, the data were segmented into multiple non-overlapping segments using a sliding window of 1s. Similar to the SEED and SEED-IV datasets, DE features were extracted for each channel of each segment at five frequency bands. Finally, the DE features from 32 channels and 5 bands were formed into a feature matrix of shape $32\times 5$ as input to the model. During the experiments, we divided the continuous labels using a fixed threshold of 5 to convert them to binary classification tasks (low / high).
A.3 Handling Different Number of Channels
Since the SEED-series datasets and the DEAP dataset contain different numbers of electrodes (channels), we require channel processing before inputting DE features into the model. Specifically, we consider the fine-tuning dataset as an anchor. When the number of channels in the fine-tuning dataset is less than in the pre-training dataset, we select data from the corresponding channels in the pre-training dataset and drop the data from the redundant channels as inputs (e.g., pre-training on SEED and fine-tuning on DEAP). Conversely, when the number of channels in the fine-tuning dataset is greater than in the pre-training dataset, we fill the missing channel data with zeros in the pre-training dataset to match the fine-tuning dataset (e.g., pre-training on DEAP and fine-tuning on SEED).
Appendix B Implementation Details
To reduce computational load while maintaining model performance, we adopt a lightweight design for each module of SCMM. Specifically, we use a 3-layer 1D CNN for the encoder $E$ and a 2-layer MLP for the projector $P$. For the lightweight decoder $D$, we utilize a single-layer MLP for reconstruction. Regarding hyperparameter selection in the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking, i.e., the percentage of random masking and channel masking is 0.9 and 0.1, respectively. We use the negative of cosine similarity as the metric function $Dist(\cdot,\cdot)$, and we set $\alpha$ to 0.5, $\tau_{w}$ to 0.05 and $\tau_{c}$ to 0.5 for soft contrastive learning. We use Adam as the optimizer with a learning rate of $5\times 10^{-4}$ and an L2-norm penalty coefficient $3\times 10^{-4}$. The pre-training process is conducted over 200 epochs with a batch size of 256. We save the model parameters $\theta$ from the final epoch as the pre-trained SCMM. In the fine-tuning stage, we input the encoded embeddings $h_{i}$ into a classifier consisting of a 2-layer fully connected network for final emotion recognition. The Adam optimizer is utilized with a learning rate of $5\times 10^{-4}$ and a weight decay of $3\times 10^{-4}$. The number of fine-tuning epochs is set to 50 for the SEED and SEED-IV datasets and 500 for the DEAP dataset, with a batch size of 128. For efficient deployment and testing of the model, the pre-trained SCMM is optimized solely using cross-entropy loss during fine-tuning. All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce 3090 GPU.
   Pre-training    Fine-tuning
Encoder 3-layer 1D CNN
Projector 2-layer MLP
Decoder single-layer MLP
Classifier - 2-layer MLP
Masking ratio $r$ 0.5 -
Threshold $\mu$ 0.1 -
Upper bound $\alpha$ 0.5 -
Sharpness $\tau_{w}$ 0.05 -
Temperature $\tau_{c}$ 0.5 -
Epoch 200 50, 500
Optimizer Adam
Learning rate $5\times 10^{-4}$
Weight decay $3\times 10^{-4}$
Batch size 256 128
Table 7: Hyperparameter settings of SCMM.
Appendix C Baseline Methods and Experimental Settings
C.1 Baseline Methods
We compare the proposed SCMM against eight competitive methods, including four traditional deep learning methods: BiDANN (Li et al. 2018), TANN (Li et al. 2021), PR-PL (Zhou et al. 2023) and E2STN (Zhou et al. 2024), as well as four self-supervised learning models: SimCLR (Chen et al. 2020; Tang et al. 2021), Mixup (Zhang et al. 2018; Wickstrøm et al. 2022), MAE (He et al. 2022) and JCFA (Liu et al. 2024a). Details of the eight baseline methods are summarized as follows.
  • BiDANN (Li et al. 2018): The bi-hemispheres domain adversarial neural network mapped the EEG data of both left and right hemispheres into discriminative feature spaces separately to address domain shifts in EEG-based emotion recognition tasks.
  • TANN (Li et al. 2021): The transferable attention neutral network is a novel transfer learning methods which learned the discriminative information from EEG signals using local and global attention mechanisms.
  • PR-PL (Zhou et al. 2023): The prototypical representation based pairwise learning framework adopted pairwise learning to model the relative relationships between EEG sample pairs in terms of prototypical representations, addressing the critical issues of individual differences and noise labels in cross-subject scenarios.
  • E2STN (Zhou et al. 2024): The emotional EEG style transfer network integrated content information from the source domain with style information from the target domain, achieving superior performance in cross-corpus EEG-based emotion recognition tasks.
  • SimCLR (Chen et al. 2020; Tang et al. 2021): A seminal work in self-supervised contrastive learning, first proposed for computer vision, has been extended to human activity recognition.
  • Mixup (Zhang et al. 2018; Wickstrøm et al. 2022): A novel CL-based data augmentation method that aimed to correctly predict the mixing proportion of two samples, has been applied to time series analysis.
  • MAE (He et al. 2022): A groundbreaking work in the field of mask modeling, which proposed to reconstruct the masked portion based on the unmasked portion of the masked sample, has achieved remarkable success in a wide range of fields.
  • JCFA (Liu et al. 2024a): The joint contrastive learning framework performed joint contrastive learning across two domains to synchronize the time- and frequency-based embeddings of the same EEG sample in the latent time-frequency space, achieving state-of-the-art performance in cross-corpus scenarios.
To ensure a fair comparison, we adopt the same encoder, projector, decoder, and classifier structures for SimCLR, Mixup and MAE as used in SCMM. We use the default hyperparameters reported in the original paper for all methods in our experiments, unless otherwise specified. Additionally, for BiDANN, TANN, PR-PL, E2STN, MAE, and SCMM, the input samples are preprocessed 1-s DE features. In contrast, SimCLR, Mixup, and JCFA use preprocessed 1-s EEG signals as inputs, in accordance with the specific design of each model.
C.2 Experimental Settings
We adopt a cross-corpus subject-independent experimental protocol in the experiments, following the setup used in JCFA. Specifically, samples from one dataset are used for pre-training, while samples of each subject from another dataset are used individually for fine-tuning and testing. During the fine-tuning process, we employ a leave-trials-out setting, where samples from a part of trials in each session of each subject in the fine-tuning dataset are used for fine-tuning, and the remaining trials are used for testing. For example, SCMM is pre-trained on the SEED-IV dataset and fine-tuned on the SEED dataset. The allocation of samples for fine-tuning/testing is as follows: 9/6 for the SEED dataset, 12/6 for the SEED-IV 3-category dataset, 16/8 for the SEED-IV 4-category dataset, and 24/16 for the DEAP dataset. This approach effectively avoids information leakage. The detailed experimental settings and data division for pre-training and fine-tuning are summarized in Table 8.
Evaluations Scenarios    Pre-training    Fine-tuning / Testing
Same-Class SEED-IV3 $\rightarrow$ SEED3 SEED-IV, 3-category SEED: 9 / 6 trials in each session per subject
SEED3 $\rightarrow$ SEED-IV3 SEED SEED-IV, 3-category: 12 / 6 trials in each session per subject
Different-Class SEED-IV4 $\rightarrow$ SEED3 SEED-IV, 4-category SEED: 9 / 6 trials in each session per subject
SEED3 $\rightarrow$ SEED-IV4 SEED SEED-IV, 4-category: 16 / 8 trials in each session per subject
DEAP $\rightarrow$ SEED3 DEAP SEED: 9 / 6 trials in each session per subject
SEED3 $\rightarrow$ DEAP (Valence) SEED DEAP (Valence): 24 / 16 trials per subject
SEED3 $\rightarrow$ DEAP (Arousal) SEED DEAP (Arousal): 24 / 16 trials per subject
Table 8: Pre-training and fine-tuning scenarios under two conditions for cross-corpus EEG-based emotion recognition.
Appendix D Masking Strategy
This paper introduces a novel hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships. To compare our approach with traditional masking strategies, we explore three different masking rules: random masking, channel masking, and hybrid masking. Figure 3 illustrates the difference between these three masking strategies.
(1) Random Masking: Generate masks using a binomial distribution to randomly mask samples in the feature dimension, setting the values of the masked features to zero.
(2) Channel Masking: Generate masks using a binomial distribution to randomly mask samples in the channel dimension, setting the values of all features within the masked channels to zero.
(3) Hybrid Masking: Generate a probability matrix using a uniform distribution that proportionally mixes masks generated from random masking and channel masking.
Our proposed hybrid masking strategy is highly flexible and can be extended to various datasets by integrating multiple masking strategies in different ratios, which is exceptionally suitable for data with rich semantic information. This approach effectively generates more diverse masked samples, encouraging the model to comprehensively capture the inherent relationships of the data.
Appendix E Hyperparameter Analysis
We conduct comprehensive experiments to verify the hyperparameter sensitivity of SCMM on the SEED and SEED-IV 3-category datasets. The hyperparameters examined include the masking ratio $r$, probability threshold $\mu$, metric function $Dist(\cdot,\cdot)$, upper bound $\alpha$, sharpness $\tau_{w}$, and temperature $\tau_{c}$. The complete experimental results are presented in the following sections.
E.1 Masking Ratio
Table 9 shows the model performance in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments when using different masking ratios (0.1, 0.25, 0.5, and 0.75). Experimental results demonstrate that the proposed SCMM consistently achieves the best performance when the masking ratio is set to $r=0.5$.
Ratio ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
0.1 91.19 / 08.07 85.05 / 11.68
0.25 91.50 / 07.69 85.28 / 09.91
0.5 91.61 / 07.56 87.24 / 08.35
0.75 90.24 / 08.08 83.49 / 13.19
Table 9: Hyperparameter sensitivity analysis of masking ratio $r$ on SEED and SEED-IV.
E.2 Probability Threshold
Table 10 presents the model performance using different probability thresholds (ranging from 0 to 1) for hybrid masking. Here, $\mu=0$ refers to the exclusive use of the random masking strategy, while $\mu=1$ indicates the exclusive use of the channel masking strategy. Experimental results show that SCMM achieves the best performance in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments when the probability threshold is set to $\mu=0.1$ (i.e., the ratio of random masking and channel masking is 9:1).
Threshold ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
0 (Random) 90.30 / 08.80 84.63 / 10.99
0.1 91.61 / 07.56 87.24 / 08.35
0.2 90.93 / 08.18 86.51 / 08.44
0.3 90.84 / 07.70 86.66 / 09.98
0.4 89.34 / 08.70 86.94 / 09.81
0.5 91.00 / 08.22 87.16 / 11.09
0.6 89.93 / 08.90 85.83 / 09.65
0.7 91.08 / 07.64 85.14 / 11.60
0.8 89.52 / 08.29 86.10 / 12.48
0.9 89.42 / 08.18 86.05 / 11.75
1 (Channel) 90.25 / 08.68 85.91 / 10.97
Table 10: Hyperparameter sensitivity analysis of threshold $\mu$ on SEED and SEED-IV.
E.3 Metric Function
We evaluate the model performance of SCMM using different metric functions in soft assignments $w(\cdot,\cdot)$. Experimental results in Table 11 indicate that SCMM performs best on the SEED and SEED-IV 3-category datasets when cosine similarity is used as the metric function.
Metrics ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
Manhattan 90.09 / 09.06 85.83 / 11.14
Euclidean 90.86 / 08.57 85.40 / 13.51
Cosine 91.61 / 07.56 87.24 / 08.35
Table 11: Hyperparameter sensitivity analysis of metric function $Dist(\cdot,\cdot)$ on SEED and SEED-IV.
E.4 Upper Bound
We assess the impact of different upper bounds on calculating soft assignments $w(\cdot,\cdot)$. To restrict the soft assignments to the range of 0 to 1, we explore the model performance of SCMM with upper bounds of 0.25, 0.5, 0.75, and 1. Experimental results in Table 12 illustrate that the proposed SCMM achieves the best performance in the SEED-IV3 $\rightarrow$ SEED3 experiment when the upper bound is set to $\alpha=0.5$.
Upper Bound ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3
0.25 91.21 / 08.54
0.5 91.61 / 07.56
0.75 91.12 / 08.25
1 90.50 / 08.00
Table 12: Hyperparameter sensitivity analysis of upper bound $\alpha$ in the SEED-IV3 $\rightarrow$ SEED3 experiment.
E.5 Sharpness
Table 13 explore the impact of different sharpness parameters in the SEED-IV3 $\rightarrow$ SEED3 experiment, as shown in Table 13. Experimental results indicate that our model achieves the best performance with a sharpness parameter of $\tau_{w}=0.05$.
Sharpness ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3
0.01 90.93 / 09.13
0.05 91.61 / 07.56
0.1 91.08 / 09.00
0.2 90.59 / 07.40
0.5 90.63 / 09.16
1 90.09 / 08.62
2 89.58 / 09.26
Table 13: Hyperparameter sensitivity analysis of sharpness $\tau_{w}$ in the SEED-IV3 $\rightarrow$ SEED3 experiment.
E.6 Temperature
We conduct comparison experiments on the SEED and SEED-IV 3-category datasets to explore the impact of temperature parameter on model performance. Experimental results in Table 14 show that SCMM consistently achieves the best performance in both experiments when the temperature parameter is set to $\tau_{c}=0.5$.
Temperature ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
0.05 90.97 / 08.17 85.95 / 11.14
0.2 91.16 / 07.85 86.82 / 11.46
0.5 91.61 / 07.56 87.24 / 08.35
2 90.62 / 08.66 86.73 / 10.18
Table 14: Hyperparameter sensitivity analysis of temperature $\tau_{c}$ on SEED and SEED-IV.
F Model Performance and Computational Complexity Analysis of SCMM
Metrics 1% of labeled data
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
ACC / STD (%) 74.98 / 17.00 71.35 / 16.29 73.20 / 16.09 60.59 / 23.72
Time Cost (s) 232.35 97.11 232.35 130.47
5% of labeled data
ACC / STD (%) 95.23 / 06.98 87.20 / 16.12 90.51 / 12.43 86.01 / 09.76
Time Cost (s) 240.48 100.23 240.48 132.90
10% of labeled data
ACC / STD (%) 99.86 / 00.31 90.04 / 16.94 99.72 / 00.80 94.43 / 05.60
Time Cost (s) 250.81 102.03 250.81 135.67
20% of labeled data
ACC / STD (%) 100.0 / 00.00 94.16 / 08.52 100.0 / 00.00 98.91 / 02.62
Time Cost (s) 267.21 105.48 267.21 139.31
Model Parameters: 1.52 M
Table 15: Model performance and computational complexity of SCMM when fine-tuning with limited labeled data (1%, 5%, 10%, and 20%) on SEED and SEED-IV.
To investigate the trade-off between model performance and computational complexity in few-shot scenarios, we assess the classification accuracy, time cost, and the number of trainable parameters of SCMM when fine-tuning with limited labeled data. Experimental results in Table 15 show that the model performance of SCMM significantly improves as the number of fine-tuning labeled samples increases. Meanwhile, the inference time cost remains low due to our lightweight design. Additionally, our model achieves superior classification performance in cross-corpus EEG-based emotion recognition tasks with very few parameters. In summary, the proposed SCMM effectively balances model performance and computational complexity.
G Visualization
G.1 Intra- and Inter-Class Similarities
Figure 7: Intra- and inter-class cosine similarities of embeddings learned by hard CL and soft CL. (a) is the average intra-class cosine similarity, and (b) is the average inter-class cosine similarity. Best viewed in color.
To assess the quality of the embeddings learned by SCMM, we randomly select one subject from the SEED dataset, and calculate both the average intra- and inter-class cosine similarities between the learned embeddings of all test samples, as shown in Fig. 7. It is evident that the proposed SCMM produces embeddings with higher intra-class similarity compared to traditional hard contrastive learning. In addition, the average inter-class similarity of the embeddings learned by SCMM is significantly lower than that of hard CL. In summary, visualization results confirm that the soft contrastive learning strategy designed in SCMM effectively clusters samples within the same category while distinctly separating samples from different categories, thus enhancing the model’s discriminative capabilities.
G.2 Reconstruction Quality
Figure 8: Comparison of reconstruction quality. We visualize the reconstruction results to compare the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM). This case is from the DEAP dataset. Best viewed in color.
To verify the effectiveness of the aggregator designed in SCMM, we compare the reconstruction quality of the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM) on the DEAP dataset. For clarity, we flatten both the original input sample and the reconstructed sample into one-dimensional vectors with dimensions $C\times F$ (channels $\times$ features). The results depicted in Fig. 8 illustrate that our model achieves lower reconstruction loss (MSE) and better sample reconstruction.

### v2

@@
 EEG-SCMM: Soft Contrastive Masked Modeling for Cross-Corpus EEG-Based Emotion Recognition
-Abstract
-Emotion recognition using electroencephalography (EEG) signals has garnered widespread attention in recent years.
-However, existing studies have struggled to develop a sufficiently generalized model suitable for different datasets without re-training (cross-corpus).
-This difficulty arises because distribution differences across datasets far exceed the intra-dataset variability.
-To solve this problem, we propose a novel Soft Contrastive Masked Modeling (SCMM) framework.
-Inspired by emotional continuity, SCMM integrates soft contrastive learning with a new hybrid masking strategy to effectively mine the ”short-term continuity” characteristics inherent in human emotions.
-During the self-supervised learning process, soft weights are assigned to sample pairs, enabling adaptive learning of similarity relationships across samples.
-Furthermore, we introduce an aggregator that weightedly aggregates complementary information from multiple close samples based on pairwise similarities among samples to enhance fine-grained feature representation, which is then used for original sample reconstruction.
-Extensive experiments on the SEED, SEED-IV and DEAP datasets show that SCMM achieves state-of-the-art (SOTA) performance, outperforming the second-best method by an average accuracy of 4.26% under two types of cross-corpus conditions (same-class and different-class) for EEG-based emotion recognition.
-Introduction
-Emotions are human attitudinal experiences and behavioral responses to objective things, closely related to health conditions and behavioral patterns (Wang, Zhang, and Tang 2024).
-Compared to speech (Singh and Goel 2022), gestures (Noroozi et al.
-2018), and facial expressions (Canal et al.
-2022), electroencephalography (EEG) offers a more direct and objective measurement of human emotions by capturing brain activity across various scalp locations (Hu et al.
-2019).
-Recently, researchers have increasingly emphasised EEG-based emotion recognition (Zhong, Wang, and Miao 2020; Zhao, Yan, and Lu 2021; Zhang, Liu, and Zhong 2022; Gao et al.
-2024), aiming to advance the development of affective brain-computer interfaces (aBCIs).
+Abstract.
+Emotion recognition using electroencephalography (EEG) signals has attracted increasing attention in recent years.
+However, existing methods often lack generalization in cross-corpus settings, where a model trained on one dataset is directly applied to another without retraining, due to differences in data distribution and recording conditions.
+To tackle the challenge of cross-corpus EEG-based emotion recognition, we propose a novel framework termed Soft Contrastive Masked Modeling (SCMM).
+Grounded in the theory of emotional continuity, SCMM integrates soft contrastive learning with a hybrid masking strategy to effectively capture emotion dynamics (refer to short-term continuity).
+Specifically, in the self-supervised learning stage, we propose a soft weighting mechanism that assigns similarity scores to sample pairs, enabling fine-grained modeling of emotional transitions and capturing the temporal continuity of human emotions.
+To further enhance representation learning, we design a similarity-aware aggregator that fuses complementary information from semantically related samples based on pairwise similarities, thereby improving feature expressiveness and reconstruction quality.
+This dual design contributes to a more discriminative and transferable representation, which is crucial for robust cross-corpus generalization.
+Extensive experiments on the SEED, SEED-IV, and DEAP datasets show that SCMM achieves state-of-the-art (SOTA) performance, outperforming the second-best method by an average accuracy of 4.26% under both same-class and different-class cross-corpus settings.
+The source code is available at https://github.com/Kyler-RL/SCMM.
+EEG; Emotion Recognition; Soft Contrastive Learning; Masked Modeling; Cross-Corpus
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755798ccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Artificial intelligence
+1. Introduction
+Emotions are human attitudinal experiences and behavioral responses to objective things, closely related to an individual’s health conditions and behavioral patterns (Wang et al., 2024).
+Compared to speech (Singh and Goel, 2022), gestures (Noroozi et al., 2018), and facial expressions (Canal et al., 2022), electroencephalography (EEG) offers a more direct and objective measurement of human emotions by capturing brain activity across various scalp locations (Hu et al., 2019).
+Therefore, researchers have increasingly emphasized EEG-based emotion recognition in recent years (Zhong et al., 2020; Zhao et al., 2021; Zhang et al., 2022, 2023), aiming to advance the development of affective brain-computer interfaces (aBCIs).
 However, three critical challenges remain to be addressed in current approaches.
-(1) Dataset Specificity.
-Most existing EEG-based emotion recognition methods are typically designed for a single dataset, necessitating model retraining when the dataset changes.
-This requirement significantly limits the model’s generalizability and scalability, hindering its application across different datasets.
-To tackle this issue, the concept of cross-corpus has been proposed, which is designed to be generalized across multiple datasets.
-A cross-corpus model is trained on one dataset and can be directly applied to another without the need for retraining from scratch.
-This concept, which originated in natural language processing (Schuller et al.
-2010; Zhang et al.
-2011), has been extended to various domains in recent years (Rayatdoost and Soleymani 2018; Chien, Yang, and Lee 2020; Ryumina, Dresvyanskiy, and Karpov 2022).
-Although existing EEG-based emotion recognition methods, such as BiDANN (Li et al.
-2018), TANN (Li et al.
-2021), and PR-PL (Zhou et al.
-2023), have achieved superior performance in within-subject or cross-subject tasks, their efficacy significantly degrades in cross-corpus scenarios, where differences in data distribution across datasets far exceed the variability within a single dataset (Rayatdoost and Soleymani 2018).
-Figure 1: (a) An illustration of emotional continuity.
+Figure 1.
+(a) An illustration of emotional continuity.
 We take the sample $x_{i}$ at second $t$ within an EEG trial as the anchor, and calculate the cosine similarity between $x_{i}$ and its neighboring sample $x_{j}$, as well as the distant sample $x_{k}$.
@@
 SCMM.
-Traditional hard contrastive learning considers the embeddings $z_{i}$ and $\widetilde{z}_{i}$ of the same sample $x_{i}$ and its augmented view $\widetilde{x}_{i}$ as positive pairs, while different samples and their augmented views are treated as negative pairs.
-When computing the contrastive loss, the weights for positives and negatives are set to 1 and 0, respectively.
-In contrast, the soft contrastive learning strategy designed in SCMM generates soft assignments for different sample pairs based on their distances in the original data space.
-(2) Data Availability.
-Current approaches for cross-domain or cross-corpus EEG-based emotion recognition rely heavily on domain adaptation techniques, which depend extensively on the availability of labeled source data and unlabeled target data.
-For example, AD-TCN (He, Zhong, and Pan 2022) learned an asymmetric mapping that adapts the target domain feature encoder to the source domain, aiming to eliminate the complicated steps of target domain labeling and improve the model performance in cross-domain scenarios.
-Similarly, E2STN (Zhou et al.
-2024) integrated content information from the source domain with style information from the target domain to create stylized emotional EEG representations.
-However, these methods require prior access to all labeled source data and unlabeled target data for model training, presenting a significant limitation due to data availability constraints (Liu et al.
-2024a).
+Traditional hard CL assigns hard values (1 and 0) to positives and negatives when computing the contrastive loss.
+In contrast, our model generates soft assignments for different sample pairs, taking into account the emotional continuity.
+[…]
+(a) An illustration of emotional continuity. (b) Hard CL vs. SCMM.
+(1) Insufficient Generalization Capability.
+Most existing EEG-based emotion recognition methods are typically designed for a single dataset, necessitating model retraining when the dataset changes.
+This requirement significantly limits the scalability and generalizability of the model, hindering its application on different datasets.
+To address this issue, the concept of cross-corpus has been proposed, which is designed to be generalized across multiple datasets.
+A cross-corpus model is trained on one dataset and can be directly applied to another without retraining from scratch.
+This concept, which originated in natural language processing (Schuller et al., 2010; Zhang et al., 2011), has been extended to various domains in recent years (Rayatdoost and Soleymani, 2018; Chien et al., 2020; Ryumina et al., 2022).
+Although existing EEG-based emotion recognition methods, such as BiDANN (Li et al., 2018b), TANN (Li et al., 2021), and PR-PL (Zhou et al., 2023), have demonstrated superior performance in within-subject or cross-subject tasks within a single dataset, their effectiveness significantly degrades in cross-corpus scenarios, where the differences in data distribution across datasets far exceed the intra-dataset variability (Rayatdoost and Soleymani, 2018).
+(2) Modeling Strategy Limitation.
+Recently, researchers have explored domain adaptation techniques to address cross-corpus EEG-based emotion recognition (He et al., 2022b; Zhou et al., 2025).
+This is motivated by their efficacy to solve the problem of domain shift (Ganin et al., 2016).
+Despite initial success, such approaches often require prior access to all labeled source data and unlabeled target data for model training.
+Considering the difficulties in collecting EEG signals as well as the time and expertise required to label them, the modeling strategy limitation of domain adaptation techniques poses a significant challenge to real-world aBCI applications.
 (3) Ignorance of Emotional Continuity.
-Unlike domain adaptation techniques, contrastive learning (CL) achieves superior performance without relying on labeled data, and has demonstrated significant potential in enhancing feature representation capabilities across various domains (Chen et al.
-2020; Radford et al.
-2021; Zhang et al.
-2022).
-Current CL-based methods for EEG-based emotion recognition consider an anchor and its augmented views as positive pairs, while treating all other samples as negatives, as shown in Fig.
+Unlike domain adaptation techniques, contrastive learning (CL) achieves superior performance without relying on labeled data, and has shown great potential in various fields (Chen et al., 2020; Radford et al., 2021; Eldele et al., 2021).
+Current CL-based methods for EEG-based emotion recognition, such as CLISA (Shen et al., 2022) and JCFA (Liu et al., 2024), consider an anchor and its augmented views as positive pairs, while treating all other samples as negatives.
+When computing the contrastive loss, the weights for positives and negatives are set to 1 and 0, respectively, as shown in Fig.
 1(b) (Hard CL).
-For example, CLISA (Shen et al.
-2022) employed contrastive learning to minimize the inter-subject differences by maximizing the similarity in EEG representations across subjects.
-JCFA (Liu et al.
-2024a) performed joint contrastive learning across three domains to align the time- and frequency-based embeddings of the same sample in the latent time-frequency space, achieving state-of-the-art (SOTA) performance in cross-corpus scenarios.
-However, psychological and neuroscientific studies have shown that emotions exhibit significant ”short-term continuity” characteristics (Davidson 1998; Houben, Van Den Noortgate, and Kuppens 2015).
-In other words, human emotions are relatively stable over certain periods, with sudden changes being rare.
+However, psychological and neuroscientific studies suggest that emotion analysis using brain signals should account for dynamic changes (Davidson, 1998; Houben et al., 2015).
+Specifically, emotions exhibit significant ”short-term continuity” characteristics, meaning that human emotions are relatively stable over certain periods, with sudden changes being rare.
 As illustrated in Fig.
 1(a), a high cosine similarity is maintained between an anchor sample $x_{i}$ and its neighboring sample $x_{j}$, and even a distant sample $x_{k}$ separated by extended periods (e.g., 60 seconds).
-Given this nature of emotions, we propose that the definition of positive pairs in CL-based EEG analysis should extend beyond just the anchor and its augmented views.
+Given this nature of emotions, we propose that the definition of positive pairs in CL-based EEG emotion analysis should extend beyond just the anchor and its augmented views.
 Instead, it should include a broader range of similar samples, especially those that are temporally proximal, as shown in Fig.
 1(b) (SCMM).
-However, existing methods like JCFA (Liu et al.
-2024a), which follow the traditional CL paradigm (Chen et al.
-2020), may incorrectly pull apart similar but not identical samples, thus failing to capture the inherent correlations of EEG signals.
-To address the aforementioned three critical issues, we propose a novel Soft Contrastive Masked Modeling (SCMM) framework for cross-corpus EEG-based emotion recognition.
+In contrast, existing methods following the traditional CL paradigm (Chen et al., 2020) may incorrectly pull apart similar but not identical samples, thus failing to capture the emotional continuity inherent in EEG signals.
+To tackle the aforementioned three critical issues, we propose a novel Soft Contrastive Masked Modeling (SCMM) framework for cross-corpus EEG-based emotion recognition.
 Unlike traditional hard CL shown in Fig.
 1(b), SCMM considers emotional continuity and incorporates soft assignments of sample pairs.
-This approach enables the model to identify the fine-grained relationships between samples in a self-supervised manner, thereby enhancing the generalizability of EEG representations.
-Comprehensive experiments on three well-recognized datasets demonstrate that SCMM consistently achieves SOTA performance, highlighting its superior capability and stability.
+This approach enables the model to identify the fine-grained relationships between different samples in a self-supervised manner, thereby enhancing the generalizability of EEG representations.
+Comprehensive experiments on three well-recognized datasets show that SCMM consistently achieves state-of-the-art (SOTA) performance, demonstrating its superior capability and stability.
 In summary, the main contributions of SCMM are outlined as follows:
-  • Inspired by the nature of emotions, we propose a novel SCMM framework to address cross-corpus EEG-based emotion recognition.
-This approach assigns soft weights to sample pairs during contrastive learning to capture the similarity relationships between different samples.
+  • We propose a novel SCMM framework to address three key challenges (insufficient generalization capability, modeling strategy limitation, and ignorance of emotional continuity) in cross-corpus EEG-based emotion recognition.
+  • Inspired by the nature of emotions, we introduce a soft weighting mechanism that assigns similarity scores to sample pairs to capture the similarity relationships between different samples.
 As a result, better feature representations of EEG signals are learned in a self-supervised manner.
-To the best of our knowledge, this is the first study to introduce soft contrastive learning into EEG-based emotion recognition.
-  • We also develop a new hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships, which is essential for enhancing contrastive learning.
-Additionally, we introduce an aggregator that weightedly aggregates complementary information from the embeddings of multiple close samples, enabling fine-grained feature learning and improving the model’s overall capability.
-  • We conduct extensive experiments on three well-known datasets (SEED, SEED-IV, and DEAP), showing that SCMM achieves consistent SOTA performance compared to eight competitive methods, surpassing the second-best method by an average accuracy of 4.26%.
-Methodology
-Figure 2: The overall framework of SCMM. The pre-training process of SCMM involves three modules: (1) hybrid masking, (2) soft contrastive learning, and (3) aggregate reconstruction.
-Problem Definition
-Given an unlabeled pre-training EEG emotion dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$ with $N$ samples, where each sample $x_{i}\in\mathbb{R}^{C\times F}$ contains $C$ channels and $F$-dimensional features, the goal is to learn a nonlinear embedding function $f_{\theta}$.
+  • We develop a new hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships, which is essential for enhancing contrastive learning.
+In addition, we introduce a similarity-aware aggregator to fuse complementary information from semantically related samples, enabling fine-grained feature learning and improving the model’s overall capability.
+  • We conduct extensive experiments on three well-known datasets (SEED, SEED-IV, and DEAP), demonstrating that SCMM achieves SOTA performance against 10 baselines, with an average accuracy improvement of 4.26% under both same-class and different-class cross-corpus settings.
+2. Related Work
+2.1. EEG-Based Emotion Recognition
+Current approaches for EEG-based emotion recognition mainly rely on two types of experimental protocols: (1) subject-dependent and (2) subject-independent.
+(1) The subject-dependent protocol trains and tests models using EEG data from the same subject within a single dataset.
+For example, Duan et al.
+(Duan et al., 2013) extracted various emotion-related features from EEG signals and used support vector machine (SVM) and k-nearest neighbors (KNN) for emotion recognition.
+Similarly, Alsolamy et al.
+(Alsolamy and Fattouh, 2016) inputted power spectral density (PSD) features into an SVM classifier to predict emotions while listening to the Quran.
+In terms of deep learning models, Zheng et al.
+(Zheng and Lu, 2015) trained a deep belief network (DBN) using differential entropy (DE) features extracted from multi-channel EEG signals for subject-dependent emotion classification.
+Additionally, Song et al.
+(Song et al., 2018) proposed a dynamical graph convolutional network (DGCNN) that dynamically learns the intrinsic relationship between different EEG channels to enhance the model’s discriminative ability.
+In general, the subject-dependent protocol tends to achieve superior performance due to its potential to introduce information leakage.
+However, this protocol fails to take into account the significant individual differences of EEG signals, thus limiting its practical applications.
+(2) The subject-independent protocol trains and tests models using EEG data from different subjects within a single dataset.
+Since transfer learning has demonstrated its potential in addressing the problem of domain shift, a series of methods have adopted it for subject-independent EEG-based emotion recognition.
+For example, Li et al.
+(Li et al., 2018b) proposed a bi-hemispheres domain adversarial neural network (BiDANN) that considers distribution shift between training and testing data and cerebral hemispheres.
+Following this, a novel transferable attention neural network (TANN) (Li et al., 2021) was introduced to learn the emotional discriminative information of EEG signals.
+To simultaneously adapt the marginal distribution and the conditional distribution, Li et al.
+(Li et al., 2019) proposed a joint distribution adaptation network (JDA) for subject-independent EEG-based emotion recognition.
+Similarly, Chen et al.
+(Chen et al., 2021) introduced a multi-source marginal distribution adaptation network (MS-MDA) to capture both domain-invariant and domain-specific features of emotional EEG signals.
+Further, Zhou et al.
+(Zhou et al., 2023) proposed a novel prototypical representation-based pairwise learning framework (PR-PL) to address individual differences and noisy labeling in emotional EEG signals.
+Despite the progress made in dealing with individual differences, these methods struggle to mitigate the distributional differences across datasets, resulting in significant performance degradation in cross-corpus scenarios.
+2.2. Cross-Corpus EEG Emotion Recognition
+To alleviate the large distributional differences across datasets, researchers have attempted to use domain adaptation techniques for cross-corpus EEG-based emotion recognition.
+For example, He et al.
+(He et al., 2022b) proposed an adversarial discriminative temporal convolutional network (AD-TCN) that integrates the adversarial discriminative learning into a temporal convolutional network for enhancing distribution matching.
+Meanwhile, Li et al.
+(Li et al., 2024) proposed a novel distillation-based domain generalization network (DBDG) to learn the discriminative and generalizable emotional features.
+Moreover, Zhou et al.
+(Zhou et al., 2025) introduced an EEG-based emotion style transfer network (E2STN) that contains the content information of the source domain and the style information of the target domain, achieving superior performance in cross-corpus scenarios.
+However, these methods require prior access to all labeled source data and unlabeled target data for model training, which is not feasible in practical applications due to the difficulties in collecting and labeling EEG data.
+Therefore, recent studies have sought to leverage contrastive learning for cross-corpus EEG-based emotion recognition.
+A noteworthy attempt is JCFA (Liu et al., 2024), which performs joint contrastive learning across three domains to align the time- and frequency-based embeddings of the same EEG sample in the latent time-frequency space, achieving SOTA performance in cross-corpus EEG-based emotion recognition tasks.
+However, such approaches following the traditional CL paradigm (Chen et al., 2020) fail to capture the emotional continuity inherent in EEG signals, resulting in relatively limited performance in cross-corpus scenarios.
+Therefore, this study aims to develop a sufficiently generalized model without relying on labeled data, which can be directly applied to different EEG emotion datasets without retraining from scratch, and achieve accurate and efficient cross-corpus generalization.
+3. Problem Formulation
+Given an unlabeled pre-training EEG emotion dataset $X=\{x_{i}\}_{i=1}^{N}$ with $N$ samples, where each sample $x_{i}\in\mathbb{R}^{C\times F}$ contains $C$ channels and $F$-dimensional features, the goal is to learn a nonlinear embedding function $f_{\theta}$.
 This function is designed to map $x_{i}$ to its representation $h_{i}$ that best describes itself by leveraging the emotional continuity inherent in EEG signals.
-Ultimately, the pre-trained model is capable of producing generalizable EEG representations that can be effectively used across different EEG emotion datasets.
-Model Architecture
+Ultimately, the pre-trained model is capable of producing generalizable EEG representations that can be effectively used across different datasets.
+Figure 2.
+The overall framework of SCMM.
+The pre-training process of SCMM involves three modules: (1) hybrid masking, (2) soft contrastive learning, and (3) aggregate reconstruction.
+Best viewed in color.
+[…]
+The overall framework of SCMM.
+4. Methodology
 The overall framework of SCMM is illustrated in Fig.
 2, which includes three main modules: hybrid masking, soft contrastive learning, and aggregate reconstruction.
-Below, we detail the specific design of each module and the pre-training process of SCMM.
-Hybrid Masking
-The selection of masking strategies is crucial for CL and masked modeling (Zhang et al.
-2024; Liu et al.
-2024b).
-For an input EEG sample $x_{i}\in\mathcal{X}$, most existing methods use random masking (Zhang, Liu, and Zhong 2022) or channel masking (Li et al.
-2022) to generate the masked sample $\widetilde{x}_{i}$.
-The random masking strategy masks samples along the feature dimension, ignoring the inter-channel relationships of EEG signals.
+Below, we will detail the specific design of each module and the pre-training process of SCMM.
+4.1. Hybrid Masking
+The selection of masking strategies is crucial for CL and masked modeling.
+For an input EEG sample $x_{i}\in X$, existing methods use random masking (Zhang et al., 2022) or channel masking (Li et al., 2022b) to generate the masked sample $\widetilde{x}_{i}$.
+The random masking strategy masks samples along the feature dimension, ignoring the inter-channel relationships of multi-channel EEG signals.
 While a large masking ratio (e.g., 75%) can mask entire portions of certain channels, it complicates the modeling process due to significant information loss.
@@
 Therefore, we develop a new hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships.
-Specifically, we first generate a random masking matrix ${\rm Mask}_{R}\in\{0,1\}$ with dimensions $C\times F$, and a channel masking matrix ${\rm Mask}_{C}\in\{0,1\}$ with dimensions $C\times F$, both derived from binomial distributions with masking ratios $r\in(0,1)$.
+Specifically, we first generate a random masking matrix ${\rm Mask}_{R}\in\{0,1\}$ with dimensions $C\times F$ and a channel masking matrix ${\rm Mask}_{C}\in\{0,1\}$ with dimensions $C\times F$, both derived from binomial distributions with the same masking ratio $r\in(0,1)$.
 Here, the element values in each row of ${\rm Mask}_{C}$ are either all 1s or all 0s.
@@
 The hybrid masking process is defined as:
-$\widetilde{x}_{i,c}=\begin{cases}x_{i,c}\odot{\rm Mask}_{R,c}&\text{if }\mu<U_{c}\leq 1\\ x_{i,c}\odot{\rm Mask}_{C,c}&\text{if }0\leq U_{c}\leq\mu\end{cases},$ (1)
-where $x_{i,c}$ represents the $c$-th channel of $x_{i}$, and $\widetilde{x}_{i,c}$ is the corresponding masked sample.
-$\odot$ denotes element-wise multiplication.
-$U_{c}$ is the probability value in the $c$-th row, and $\mu$ is a probability threshold that controls the weights of the two masking strategies.
-By integrating the hybrid masking strategy in SCMM, we enhance the diversity of masked samples, encouraging the model to learn richer and more robust feature representations that account for both channel and feature relationships within EEG signals.
-Figure 3 illustrates the differences between three masking strategies.
-Figure 3: Illustrations of different masking strategies.
-(a) compares various masking strategies, and (b) presents examples of generated masked samples using three strategies.
-The masking ratio and threshold are set to $r=0.5$ and $\mu=0.1$, respectively.
-Soft Contrastive Learning
-Traditional hard CL treats augmented views generated from the same sample as positive pairs, and those from different samples as negative pairs (Chen et al.
-2020; Eldele et al.
-2021; Yue et al.
-2022).
+(1) $\widetilde{x}_{i,c}=\begin{cases}x_{i,c}\odot{\rm Mask}_{R,c}&\text{if }\mu<U_{c}\leq 1\\ x_{i,c}\odot{\rm Mask}_{C,c}&\text{if }0\leq U_{c}\leq\mu\end{cases},$
+where $\odot$ denotes element-wise multiplication.
+$x_{i,c}$ represents the DE features of the $c$-th channel of $x_{i}$, and $\widetilde{x}_{i,c}$ is the corresponding masked sample.
+$U_{c}$ represents the probability value of $U$ in the $c$-th row, and $\mu$ is a probability threshold that controls the weights of the two masking strategies.
+By integrating the hybrid masking strategy in SCMM, we enhance the diversity of masked samples, enabling the model to learn richer feature representations that account for both channel and feature relationships of EEG signals.
+Figure 3 illustrates the differences between the three masking strategies.
+More details are presented in Appendix D.
+Figure 3. Comparison of different masking strategies. The masking ratio and threshold are set to $r=0.5$ and $\mu=0.1$, respectively. Best viewed in color. Zoom in for a better view.
+[…]
+Comparison of different masking strategies.
+4.2. Soft Contrastive Learning
+Traditional hard CL treats the same sample and its augmented views as positive pairs, while treating all other samples as negatives (Chen et al., 2020).
 During the computation of the contrastive loss, hard values (1 or 0) are assigned to sample pairs, as illustrated in Fig.
@@
 Specifically, for a given pair of samples ($x_{i}$, $x_{j}$), we first calculate the normalized distance $D(x_{i},x_{j})$ between $x_{i}$ and $x_{j}$ in the original data space as:
-$D(x_{i},x_{j})=Norm(Dist(x_{i},x_{j}))\in[0,1],$ (2)
+(2) $D(x_{i},x_{j})=Norm(Dist(x_{i},x_{j}))\in[0,1],$
 where $Dist(\cdot,\cdot)$ is a metric function used to measure the distance between sample pairs, and $Norm(\cdot)$ denotes min-max normalization.
-In the experiments, we take the negative of cosine similarity as the metric function.
 Based on the normalized distance $D(x_{i},x_{j})$, we then define a soft assignment $w{(x_{i},x_{j})}$ for each pair of samples ($x_{i}$, $x_{j}$) using the sigmoid function $\sigma(x)=1/(1+{\rm exp}(-x))$:
-$w{(x_{i},x_{j})}=2\alpha\cdot\sigma(-D(x_{i},x_{j})/\tau_{w}),$ (3)
+(3) $w{(x_{i},x_{j})}=2\alpha\cdot\sigma(-D(x_{i},x_{j})/\tau_{\rm s}),$
 where $\alpha\in[0,1]$ is a boundary parameter that controls the upper bound of soft assignments.
-$\tau_{w}$ is a sharpness parameter, where smaller values of $\tau_{w}$ result in greater differences in $w(\cdot,\cdot)$ between sample pairs, and vice versa.
-Figure 4 depicts heat maps of soft assignments $w(\cdot,\cdot)$ under different sharpness parameters $\tau_{w}$.
-Figure 4: Soft assignments $w(\cdot,\cdot)$ with different $\tau_{w}$.
-We use heat maps to visualize the soft assignments with different sharpness parameters $\tau_{w}$.
-For clarity, we set the upper bound $\alpha=1$ in the figure.
-Best viewed in color.
+$\tau_{\rm s}$ is a sharpness parameter, where smaller values of $\tau_{\rm s}$ result in greater differences in $w(\cdot,\cdot)$ between sample pairs, and vice versa.
+Figure 4 illustrates the differences between soft assignments $w(\cdot,\cdot)$ with different sharpness $\tau_{\rm s}$.
+Figure 4. Heatmaps of soft assignments $w(\cdot,\cdot)$ with different sharpness $\tau_{\rm s}$. Best viewed in color. Zoom in for a better view.
+[…]
+Soft assignments with different sharpness parameters.
 Leveraging the soft assignments for all sample pairs, we propose a soft contrastive loss to refine the traditional hard contrastive loss.
 Specifically, for a pair of projected embeddings $(z_{i},\widetilde{z}_{i})$, we first calculate the softmax probability of the relative similarity among all similarities as:
-$p(z_{i},\widetilde{z}_{i})=\frac{{\rm exp}({\rm sim}(z_{i},\widetilde{z}_{i})/\tau_{c})}{\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{c})},$ (4)
-where ${\rm sim(\cdot,\cdot)}$ refers to the cosine similarity, and $\tau_{c}$ is a temperature parameter used to adjust the scale.
+(4) $p(z_{i},\widetilde{z}_{i})=\frac{{\rm exp}({\rm sim}(z_{i},\widetilde{z}_{i})/\tau_{\rm c})}{\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{\rm c})},$
+where ${\rm sim(\cdot,\cdot)}$ refers to the cosine similarity, and $\tau_{\rm c}$ is a temperature parameter used to adjust the scale.
 Based on $p(z_{i},\widetilde{z}_{i})$, the soft contrastive loss is then defined as:
-$\begin{split}\mathcal{L}_{C,i}&=-{\rm log}\thinspace p(z_{i},\widetilde{z}_{i})\\ &-\sum_{\begin{subarray}{c}x^{\prime}\in\Omega\backslash\{x_{i},\widetilde{x}_{i}\}\\ z^{\prime}\in\mathcal{Z}\backslash\{z_{i},\widetilde{z}_{i}\}\end{subarray}}w{(x_{i},x^{\prime})}\cdot{\rm log}\thinspace p(z_{i},z^{\prime}),\end{split}$ (5)
-where $\Omega=\mathcal{X}\cup\widetilde{\mathcal{X}}$ represents the union of the data spaces of the original and masked samples.
+(5) $\begin{split}\mathcal{L}_{{\rm C},i}&=-{\rm log}\thinspace p(z_{i},\widetilde{z}_{i})\\ &-\sum_{\begin{subarray}{c}{\bm{x}}\in\mathcal{X}\backslash\{x_{i},\widetilde{x}_{i}\}\\ {\bm{z}}\in\mathcal{Z}\backslash\{z_{i},\widetilde{z}_{i}\}\end{subarray}}w{(x_{i},{\bm{x}})}\cdot{\rm log}\thinspace p(z_{i},{\bm{z}}),\end{split}$
+where $\mathcal{X}=X\cup\widetilde{X}$ represents the union of the data spaces of the original and masked samples.
 By assigning soft weights to different sample pairs, the model is encouraged to better capture the inherent correlations across different samples.
-During the training process, the final soft contrastive loss $\mathcal{L}_{C}$ is computed by summing and averaging $\mathcal{L}_{C,i}$ across all samples within a mini-batch.
-Notably, when $\forall{w(x_{i},x^{\prime})}=0$, the soft contrastive loss reduces to the traditional hard contrastive loss.
-Aggregate Reconstruction
+During the training process, the final soft contrastive loss $\mathcal{L}_{\rm C}$ is computed by summing and averaging $\mathcal{L}_{{\rm C},i}$ across all samples within a mini-batch.
+Notably, when $\forall{w(x_{i},{\bm{x}})}=0$, the soft contrastive loss reduces to the traditional hard contrastive loss.
+4.3. Aggregate Reconstruction
 To further capture the fine-grained relationships between different samples, we incorporate an aggregator for weighted aggregation and reconstruction.
-Current approaches for masked EEG modeling typically reconstruct the masked portion based on the unmasked portion of a single masked sample (Lan et al.
-2024; Pang et al.
-2024), following the learning paradigm of MAE (He et al.
-2022).
+Current approaches for masked EEG modeling typically reconstruct the masked portion based on the unmasked portion of a single masked sample (Lan et al., 2024; Pang et al., 2024), following the learning paradigm of MAE (He et al., 2022a).
 However, this single-sample reconstruction strategy overlooks the interactions between samples, leading to a complex and inadequate reconstruction process.
-To overcome this limitation, we introduce an aggregator that improves the traditional single-sample reconstruction process.
+To overcome this limitation, we introduce a similarity-aware aggregator that improves the traditional single-sample reconstruction process.
 Specifically, we first calculate the cosine similarity between each pair of projected embeddings $(z_{i},z_{j})$ within a mini-batch, resulting in a similarity matrix $\mathcal{S}$.
-Based on the pairwise similarities in $\mathcal{S}$, the aggregator then performs weighted aggregation of the embedding $h_{i}$.
-The weighted aggregation process is defined as:
-$h_{i}^{r}=\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}(\frac{{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{c})}{\sum_{z^{\prime\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime\prime})/\tau_{c})}\cdot h^{\prime}),$ (6)
+Based on the pairwise similarities in $\mathcal{S}$, the aggregator then performs weighted aggregation of the embedding $h_{i}$, defined as:
+(6) $h_{i}^{r}=\sum_{z^{\prime}\in\mathcal{Z}\backslash\{z_{i}\}}(\frac{{\rm exp}({\rm sim}(z_{i},z^{\prime})/\tau_{\rm c})}{\sum_{z^{\prime\prime}\in\mathcal{Z}\backslash\{z_{i}\}}{\rm exp}({\rm sim}(z_{i},z^{\prime\prime})/\tau_{\rm c})}\cdot h^{\prime}),$
 where $h^{\prime}\in\mathcal{H}\backslash\{h_{i}\}$ represents the encoded embedding corresponding to the projected embedding $z^{\prime}$, and $\mathcal{H}$ denotes the embedding space of the encoder $E$.
-This approach allows for a more comprehensive reconstruction by aggregating complementary information and incorporating similar features from different samples during the reconstruction process, while suppressing interference from irrelevant noise samples.
+This approach allows for a more comprehensive reconstruction by aggregating complementary information and incorporating similar features from different samples during the reconstruction process.
 Finally, the reconstructed embedding $h_{i}^{r}$ is fed into a lightweight decoder $D$ to obtain the reconstructed sample $x_{i}^{r}$.
-Following the masked modeling paradigm, we use Mean Square Error (MSE) as the reconstruction loss for model optimization, which is defined as:
-$\mathcal{L}_{R,i}=\|x_{i}-x_{i}^{r}\|_{2}^{2}.$ (7)
-Similar to the soft contrastive loss $\mathcal{L}_{C}$, the final reconstruction loss $\mathcal{L}_{R}$ is computed by summing and averaging $\mathcal{L}_{R,i}$ across all samples within a mini-batch.
-The pre-training process of SCMM
-During the pre-training process, SCMM is trained by jointly optimizing $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$. The overall pre-training loss is defined as:
-$\mathcal{L}=\lambda_{C}\mathcal{L}_{C}+\lambda_{R}\mathcal{L}_{R},$ (8)
-where $\lambda_{C}$ and $\lambda_{R}$ are trade-off hyperparameters, which are adaptively adjusted according to the homoscedastic uncertainty of each loss item.
-Algorithm 1 details the pre-training process of the proposed SCMM.
+Following the masked modeling paradigm, we use Mean Squared Error (MSE) as the reconstruction loss for model optimization, which is defined as:
+(7) $\mathcal{L}_{{\rm R},i}=\|x_{i}-x_{i}^{r}\|_{2}^{2}.$
+Similar to the soft contrastive loss $\mathcal{L}_{\rm C}$, the final reconstruction loss $\mathcal{L}_{\rm R}$ is computed by summing and averaging $\mathcal{L}_{{\rm R},i}$ across all samples within a mini-batch.
+[…]
+Methods Same-Class Different-Class
+SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
+Transfer Learning
+DANN (Ganin et al., 2016)* 51.91 / 09.27 45.90 / 03.60 - -
+BiDANN (Li et al., 2018b) 49.24 / 10.49 60.46 / 11.17 - -
+TANN (Li et al., 2021) 58.41 / 07.16 60.75 / 10.61 - -
+PR-PL (Zhou et al., 2023)* 61.01 / 10.55 58.74 / 10.71 - -
+E2STN (Zhou et al., 2025) 60.51 / 05.41 61.24 / 15.14 - -
+Self-Supervised Learning
+SimCLR (Chen et al., 2020; Tang et al., 2020)* 47.27 / 08.44 46.89 / 13.41 44.19 / 09.28 42.03 / 10.05
+Mixup (Zhang, 2017; Wickstrøm et al., 2022)* 56.86 / 16.83 55.70 / 16.28 54.55 / 17.95 45.79 / 15.16
+TS-TCC (Eldele et al., 2021)* 55.38 / 11.65 49.43 / 09.44 52.30 / 12.49 44.57 / 06.08
+MAE (He et al., 2022a)* 86.49 / 10.57 83.87 / 08.53 86.02 / 08.96 76.74 / 09.18
+JCFA (Liu et al., 2024) 67.53 / 12.36 62.40 / 07.54 65.99 / 14.04 52.67 / 05.86
+SCMM (Ours) 91.61 / 07.56 (+05.12) 87.24 / 08.35 (+03.37) 91.26 / 07.91 (+05.24) 80.89 / 08.69 (+04.15)
+4.4. The Pre-training Process of SCMM
+During the pre-training process, SCMM is trained by jointly optimizing $\mathcal{L}_{\rm C}$ and $\mathcal{L}_{\rm R}$. The overall pre-training loss is defined as:
+(8) $\mathcal{L}_{\rm pret}=\lambda_{\rm C}\mathcal{L}_{\rm C}+\lambda_{\rm R}\mathcal{L}_{\rm R},$
+where $\lambda_{\rm C}$ and $\lambda_{\rm R}$ are trade-off hyperparameters that are adaptively adjusted according to the homoscedastic uncertainty of each loss item (Kendall et al., 2018).
+Algorithm 1 details the pre-training process of SCMM.
 Algorithm 1 The pre-training process of SCMM.
-Input: Unlabeled pre-training dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$. The number of pre-training epochs.
-The pre-training process:
+0:
+0: - Unlabeled pre-training EEG emotion dataset $X=\{x_{i}\}_{i=1}^{N}$. The number of pre-training epochs.
+0:
 1: Randomly initialize the model parameters $\theta$;
 2:for $epoch=1$ to epochs do
-2:      /*All operations are performed within a mini-batch*/
+2:   // All operations are performed within a mini-batch
 3:  Generate the masked sample $\widetilde{x}_{i}$ of each input EEG sample $x_{i}$ using hybrid masking in Eq. (1);
@@
 7:  Generate the soft assignment $w(x_{i},x_{j})$ for each pair of samples $(x_{i},x_{j})$ using Eq. (3);
-8:     Compute the soft contrastive loss $\mathcal{L}_{C}$ using Eq. (5);
+8:  Compute the soft contrastive loss $\mathcal{L}_{\rm C}$ using Eq. (5);
 9:  Compute the pairwise cosine similarity for each pair of projected embeddings $(z_{i},z_{j})$;
@@
 11:  Reconstruct $x_{i}^{r}$ by feeding $h_{i}^{r}$ into $D$;
-12:     Compute the reconstruction loss $\mathcal{L}_{R}$ using Eq. (7);
-13:     Compute the pre-training loss $\mathcal{L}$ using Eq. (8);
+12:  Compute the reconstruction loss $\mathcal{L}_{\rm R}$ using Eq. (7);
+13:  Compute the pre-training loss $\mathcal{L}_{\rm pret}$ using Eq. (8);
 14:  Update the model parameters $\theta$;
@@
 16:return  The pre-trained SCMM model $f_{\theta}$.
-Experiments
+5. Experiments
+5.1. Datasets
+We conduct extensive experiments on three public datasets, SEED (Zheng and Lu, 2015), SEED-IV (Zheng et al., 2018), and DEAP (Koelstra et al., 2011), to evaluate the model performance of SCMM in cross-corpus EEG-based emotion recognition tasks.
+These datasets are diverse in terms of EEG equipment, emotional stimuli, data specifications, labeling approaches, and subjects, making them well-suited for assessing the model’s efficacy in cross-corpus scenarios.
+In the experiments, we use 1-s (SEED and DEAP) and 4-s (SEED-IV) DE features as inputs, respectively.
+Detailed descriptions of the datasets and pre-processing procedures are provided in Appendix A.
+5.2. Implementation Details
+In the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking.
+We use the negative of cosine similarity as $D(\cdot,\cdot)$, and we set $\alpha$ to 0.5, $\tau_{\rm s}$ to 0.05, and $\tau_{\rm c}$ to 0.5 for soft CL.
+We use Adam as optimizer with an initial learning rate of 5e-4 and a weight decay of 3e-4.
+The pre-training process is conducted over 200 epochs with a batch size of 256.
+We save the model parameters $\theta$ from the final epoch as the pre-trained SCMM.
+In the fine-tuning stage, we input the encoded embedding $h_{i}$ into an emotion classifier consisting of a 2-layer fully connected network for final emotion recognition.
+For efficient deployment and testing of the model, the pre-trained SCMM is optimized solely using cross-entropy loss during fine-tuning.
+The fine-tuning process is conducted over 50 epochs with a batch size of 128.
+All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce RTX 3090 GPU.
+We release the source code of SCMM at https://github.com/Kyler-RL/SCMM.
+Further implementation details can be found in Appendix B.
+5.3. Baseline Models and Experimental Settings
+[…]
+Table 2. Experimental results on SEED and DEAP under the different-class cross-corpus condition.
+Methods DEAP $\rightarrow$ SEED3 SEED3 $\rightarrow$ DEAP (Valence) SEED3 $\rightarrow$ DEAP (Arousal)
+SimCLR (Chen et al., 2020; Tang et al., 2020)* 53.12 / 13.12 53.75 / 03.61 51.79 / 04.54
+Mixup (Zhang, 2017; Wickstrøm et al., 2022)* 48.75 / 14.37 60.62 / 08.68 60.11 / 07.69
+TS-TCC (Eldele et al., 2021)* 49.37 / 12.50 56.25 / 03.46 54.13 / 04.45
+MAE (He et al., 2022a)* 83.69 / 10.10 72.19 / 07.24 70.50 / 06.30
+JCFA (Liu et al., 2024) 64.69 / 12.28 61.59 / 06.26 61.06 / 07.37
+SCMM (Ours) 91.70 / 08.07 (+08.01) 73.96 / 06.75 (+01.77) 72.66 / 05.67 (+02.16)
+Table 3. Ablation study on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
 Methods Same-Class Different-Class
 SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
-BiDANN (Li et al. 2018) 49.24 / 10.49 60.46 / 11.17 - -
-TANN (Li et al. 2021) 58.41 / 07.16 60.75 / 10.61 - -
-PR-PL (Zhou et al. 2023)* 61.01 / 10.55 58.74 / 10.71 - -
-E2STN (Zhou et al. 2024) 60.51 / 05.41 61.24 / 15.14 - -
-SimCLR (Tang et al. 2021)* 47.27 / 08.44 46.89 / 13.41 44.19 / 09.28 42.03 / 10.05
-Mixup (Wickstrøm et al. 2022)* 56.86 / 16.83 55.70 / 16.28 54.55 / 17.95 45.79 / 15.16
-MAE (He et al. 2022)* 86.49 / 10.57 83.87 / 08.53 86.02 / 08.96 76.74 / 09.18
-JCFA (Liu et al. 2024a)* 67.53 / 12.36 62.40 / 07.54 65.99 / 14.04 52.67 / 05.86
-SCMM (Ours) 91.61 / 07.56 (+05.12) 87.24 / 08.35 (+03.37) 91.26 / 07.91 (+05.24) 80.89 / 08.69 (+04.15)
-Table 1: Experimental results on SEED and SEED-IV under two cross-corpus conditions: same-class and different-class.
-”*” indicates that the results are reproduced by ourselves.
-A $\rightarrow$ B denotes that A is the pre-training dataset, while B is the dataset for model fine-tuning and testing.
-Best results are highlighted in bold, while the second-best results are underlined.
-Hard $\mathcal{L}_{C}$ Soft $\mathcal{L}_{C}$ $\mathcal{L}_{R}$ Same-Class Different-Class
-SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
-90.08 / 09.24 84.43 / 11.83 90.26 / 08.71 77.79 / 08.04
-90.73 / 08.48 85.07 / 11.05 90.96 / 08.36 78.32 / 07.19
-89.68 / 09.32 84.24 / 11.90 89.45 / 09.10 77.24 / 09.14
-90.30 / 07.94 85.95 / 08.74 90.91 / 08.61 79.82 / 07.00
-91.61 / 07.56 87.24 / 08.35 91.26 / 07.91 80.89 / 08.69
-Table 2: Ablation study on SEED and SEED-IV under two cross-corpus conditions: same-class and different-class.
-Datasets
-We conduct extensive experiments on three publicly available datasets, SEED (Zheng and Lu 2015), SEED-IV (Zheng et al.
-2018), and DEAP (Koelstra et al.
-2011), to evaluate the model performance of SCMM.
-These datasets are diverse in terms of EEG equipment, emotional stimuli, data specifications, labeling approaches and subjects, making them well-suited for assessing the model’s efficacy in cross-corpus EEG-based emotion recognition tasks.
-In our experiments, we use differential entropy (DE) features as inputs.
-Detailed descriptions of the datasets and pre-processing procedures are provided in Appendix A.
-Implementation Details
-In the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking.
-We set $\alpha$ to 0.5, $\tau_{w}$ to 0.05, and $\tau_{c}$ to 0.5 for soft CL.
-The Adam optimizer is utilized with a learning rate of 5e-4 and a weight decay of 3e-4.
-The pre-training process is conducted over 200 epochs with a batch size of 256.
-We save the parameters $\theta$ from the final epoch as the pre-trained model.
-In the fine-tuning stage, we input the embeddings $h_{i}$ generated by the encoder $E$ into an emotion classifier consisting of a 2-layer MLP for final emotion recognition.
-The classifier is optimized using cross-entropy (CE) loss, and the fine-tuning process is conducted over 50 epochs with a batch size of 128.
-All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce RTX 3090 GPU.
-Further implementation details and hyperparameter analysis can be found in Appendix B and Appendix E.
-Baselines and Experimental Setup
-We compare the proposed SCMM against eight competitive baseline methods, including four conventional deep learning methods: BiDANN (Li et al.
-2018), TANN (Li et al.
-2021), PR-PL (Zhou et al.
-2023), and E2STN (Zhou et al.
-2024), as well as four self-supervised learning models: SimCLR (Chen et al.
-2020; Tang et al.
-2021), Mixup (Zhang et al.
-2018; Wickstrøm et al.
-2022), MAE (He et al.
-2022), and JCFA (Liu et al.
-2024a).
-Notably, E2STN and JCFA are specifically designed for cross-corpus EEG-based emotion recognition.
-To ensure a fair comparison, we adopt a cross-corpus subject-independent protocol in our experiments, consistent with the setup used by JCFA.
-We evaluate the model performance using the average accuracy and standard deviation (ACC / STD %) across all subjects in the test set.
-More details about baseline methods and experimental settings are provided in Appendix C.
-Experimental Results
-Methods DEAP $\rightarrow$ SEED3 SEED3 $\rightarrow$ DEAP (Valence) SEED3 $\rightarrow$ DEAP (Arousal)
-SimCLR (Tang et al. 2021)* 53.12 / 13.12 53.75 / 03.61 51.79 / 04.54
-Mixup (Wickstrøm et al. 2022)* 48.75 / 14.37 60.62 / 08.68 60.11 / 07.69
-MAE (He et al. 2022)* 83.69 / 10.10 72.19 / 07.24 70.50 / 06.30
-JCFA (Liu et al. 2024a)* 64.69 / 12.28 61.59 / 06.26 61.06 / 07.37
-SCMM (Ours) 91.70 / 08.07 (+08.01) 73.96 / 06.75 (+01.77) 72.66 / 05.67 (+02.16)
-Table 3: Experimental results for cross-corpus EEG-based emotion recognition on SEED and DEAP.
-Evaluation on the SEED and SEED-IV datasets
+w/o $\mathcal{L}_{\rm C}$ 89.68 / 09.32 84.24 / 11.90 89.45 / 09.10 77.24 / 09.14
+w/o $\mathcal{L}_{\rm R}$ 90.73 / 08.48 85.07 / 11.05 90.96 / 08.36 78.32 / 07.19
+SCMM 91.61 / 07.56 87.24 / 08.35 91.26 / 07.91 80.89 / 08.69
+5.4. Results Analysis and Comparison
 To fully validate the model performance of SCMM, we conduct extensive experiments under two cross-corpus conditions: (1) same-class and (2) different-class.
-(1) For the same-class cross-corpus validation, we conduct two experiments on the SEED and SEED-IV 3-category datasets: pre-training on SEED-IV and fine-tuning on SEED (SEED-IV3 $\rightarrow$ SEED3), and pre-training on SEED and fine-tuning on SEED-IV (SEED3 $\rightarrow$ SEED-IV3).
+Appendix E presents the within-dataset validation experimental results of SCMM on SEED and SEED-IV.
+Full results are provided in Appendix F.
+(1) Same-Class.
+We first conduct two experiments on the SEED and SEED-IV 3-category datasets: pre-training on SEED-IV and fine-tuning on SEED (SEED-IV3 $\rightarrow$ SEED3), and pre-training on SEED and fine-tuning on SEED-IV (SEED3 $\rightarrow$ SEED-IV3).
 In both experiments, all samples corresponding to fear emotions in the SEED-IV dataset are excluded.
@@
 Specifically, our model achieves classification accuracies of 91.61% and 87.24% with standard deviations of 7.56% and 8.35% in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments, outperforming the second-best method MAE by accuracies of 5.12% and 3.37%, respectively.
-(2) For the different-class cross-corpus validation, we conduct another two experiments, denoted as SEED-IV4 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV4.
+In addition, the proposed SCMM is significantly better than transfer learning methods, highlighting its superiority.
+(2) Different-Class.
+We then conduct two experiments on the SEED and SEED-IV 4-category datasets, denoted as SEED-IV4 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV4.
 These experiments aim to evaluate the model performance when the pre-training and fine-tuning datasets contain different emotion categories.
+In the experiments, we exclude transfer learning methods since they are not suitable for scenarios where the training and testing datasets contain different emotion categories.
 Experimental results in the right two columns of Table 1 demonstrate that SCMM achieves the best performance in both experiments.
 Specifically, our model achieves classification accuracies of 91.26% and 80.89% with standard deviations of 7.91% and 8.69% in the SEED-IV4 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV4 experiments, surpassing the second-best method MAE by 5.24% and 4.15% in accuracies, respectively.
-Note that the traditional CL-based models SimCLR, Mixup, and JCFA exhibit relatively poor performance, primarily due to their use of raw EEG signals as inputs.
-In addition, SCMM significantly outperforms the traditional masked modeling framework MAE, highlighting the superiority of SCMM.
-In summary, extensive experimental results on the SEED and SEED-IV datasets confirm that our model exhibits superior performance in multiple cross-corpus EEG-based emotion recognition tasks.
-Ablation study
-To assess the validity of each module in SCMM, we conduct a comprehensive ablation study on the SEED and SEED-IV datasets.
-Specifically, we design five different models below.
-(1) Hard $\mathcal{L}_{C}$ trains the model with only hard contrastive loss.
-(2) Soft $\mathcal{L}_{C}$ trains the model with only soft contrastive loss.
-(3) $\mathcal{L}_{R}$ trains the model with only reconstruction loss.
-(4) Hard $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$ trains the model using both hard contrastive loss and reconstruction loss.
-(5) Soft $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$ trains the model using both soft contrastive loss and reconstruction loss, representing the full model.
-Table 2 presents the results of ablation experiments.
-Specifically, (1) Hard $\mathcal{L}_{C}$ and (2) Soft $\mathcal{L}_{C}$: the first two models compare the performance of hard CL and soft CL.
-Experimental results indicate that our well-designed soft contrastive learning effectively improves the classification accuracy.
-(3) $\mathcal{L}_{R}$: the third model removes the soft contrastive loss and trains the model with only reconstruction loss.
-The results show that the model performs worst in the absence of the contrastive learning constraint.
-(4) Hard $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$ and (5) Soft $\mathcal{L}_{C}$ and $\mathcal{L}_{R}$: the last two models compare the performance when combining hard and soft contrastive loss with reconstruction loss, respectively.
-Experimental results demonstrate that the full model achieves the best performance in all experiments, indicating that SCMM significantly enhances the model performance and stability by combining soft contrastive learning and aggregate reconstruction.
-This improvement is evident under different cross-corpus conditions, demonstrating the feasibility of extending SCMM to real-life aBCI applications.
-Discussion
-Method ACC / STD (%)
-SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
-OS 91.61 / 07.56 87.24 / 08.35 91.25 / 07.91 80.89 / 08.69
-ES 89.99 / 10.25 85.75 / 14.00 90.31 / 08.59 79.04 / 06.95
-DEAP $\rightarrow$ SEED3 SEED3 $\rightarrow$ DEAP (Valence) SEED3 $\rightarrow$ DEAP (Arousal) -
-OS 91.70 / 08.01 73.96 / 06.75 72.66 / 05.67 -
-ES 90.64 / 07.97 72.75 / 07.06 71.58 / 05.72 -
-Table 4: Comparison of soft contrastive learning in the original data space (OS) and embedding space (ES).
-Generalization capability analysis
-To further validate the generalization capability of SCMM, we conduct additional experiments on the SEED and DEAP datasets under a different-class cross-corpus scenario, denoted as DEAP $\rightarrow$ SEED3, SEED3 $\rightarrow$ DEAP (Valence) and SEED3 $\rightarrow$ DEAP (Arousal).
-It is notable that the EEG acquisition equipment, emotional stimuli, data specifications, labeling approaches and subjects are completely different between the two datasets.
+Additionally, the traditional CL-based models SimCLR, Mixup, TS-TCC, and JCFA exhibit relatively poor performance across all experiments due to their use of raw EEG signals as inputs.
+To further validate the generalization capability of SCMM, we conduct additional experiments on the SEED and DEAP datasets, denoted as DEAP $\rightarrow$ SEED3, SEED3 $\rightarrow$ DEAP (Valence), and SEED3 $\rightarrow$ DEAP (Arousal).
+Note that the EEG acquisition equipment, emotional stimuli, data specifications, labeling approaches, and subjects are completely different between the two datasets.
 Table 2 presents the experimental results of SCMM compared to existing methods.
 Specifically, for the DEAP $\rightarrow$ SEED3 experiment, SCMM achieves an accuracy of 91.70% with a standard deviation of 8.07%, outperforming the second-best method MAE by an accuracy of 8.01%.
-For the SEED3 $\rightarrow$ DEAP (Valence) and SEED3 $\rightarrow$ DEAP (Arousal) experiments, our model achieves classification accuracies of 73.96% and 72.66% with standard deviations of 6.75% and 5.67%, surpassing the second-best method MAE by 1.77% and 2.16% in accuracies.
-Experimental results demonstrate that the proposed SCMM maintains excellent performance even when the pre-training and fine-tuning datasets are completely different, highlighting its superior generalization capability.
-Model performance with limited fine-tuning data
-Figure 5: Model performance with limited labeled data for fine-tuning on SEED and SEED-IV.
-We further explore the model performance of SCMM on the SEED and SEED-IV datasets when fine-tuning labeled data is limited.
-Specifically, we randomly select 1%, 5%, 10% and 20% of labeled samples from the fine-tuning dataset for model fine-tuning, while the remaining samples are used for testing.
+For the SEED3 $\rightarrow$ DEAP (Valence) and SEED3 $\rightarrow$ DEAP (Arousal) experiments, SCMM achieves classification accuracies of 73.96% and 72.66% with standard deviations of 6.75% and 5.67%, surpassing the second-best method MAE by 1.77% and 2.16% in accuracy.
+The results show that our model maintains excellent performance even when the pre-training and fine-tuning datasets are completely different, highlighting its superior generalization capability.
+Further, the comparative analysis of Table 1 and Table 2 reveals that the model performance of SCMM on the same fine-tuning dataset remains stable when pre-training on different datasets.
+This suggests that our model effectively captures generalizable emotional EEG representations that are robust to dataset variations.
+In summary, extensive experimental results on SEED, SEED-IV, and DEAP confirm that our model exhibits superior performance and stability in cross-corpus EEG-based emotion recognition tasks under both same-class and different-class conditions.
+Table 4. Comparison of hard CL and soft CL on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
+Methods ACC / STD (%)
+SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
+Hard CL 90.30 / 07.94 85.95 / 08.74 90.91 / 08.61 79.82 / 07.00
+Soft CL 91.61 / 07.56 87.24 / 08.35 91.25 / 07.91 80.89 / 08.69
+DEAP $\rightarrow$ SEED3 SEED3 $\rightarrow$ DEAP (Valence) SEED3 $\rightarrow$ DEAP (Arousal) -
+Hard CL 90.87 / 08.58 73.10 / 07.34 71.98 / 06.02 -
+Soft CL 91.70 / 08.01 73.96 / 06.75 72.66 / 05.67 -
+Table 5. Comparison of soft CL in the embedding space (ES) and original data space (OS) on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
+Methods ACC / STD (%)
+SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
+ES 89.99 / 10.25 85.75 / 14.00 90.31 / 08.59 79.04 / 06.95
+OS 91.61 / 07.56 87.24 / 08.35 91.25 / 07.91 80.89 / 08.69
+DEAP $\rightarrow$ SEED3 SEED3 $\rightarrow$ DEAP (Valence) SEED3 $\rightarrow$ DEAP (Arousal) -
+ES 90.64 / 07.97 72.75 / 07.06 71.58 / 05.72 -
+OS 91.70 / 08.01 73.96 / 06.75 72.66 / 05.67 -
+6. Discussions
+6.1. Ablation Study
+To assess the validity of each module in SCMM, we conduct a comprehensive ablation study on the SEED and SEED-IV datasets.
+Table 3 presents the results of ablation experiments.
+Specifically, we design three different models below.
+(1) w/o $\mathcal{L}_{\rm C}$: the first configuration removes the soft contrastive loss and trains the model using only the reconstruction loss.
+The results show that the model performs the worst without the contrastive learning constraint.
+However, it still outperforms the MAE model based on the single-sample reconstruction paradigm.
+This suggests that our aggregate reconstruction strategy effectively improves the model performance by capturing fine-grained inter-sample relationships.
+(2) w/o $\mathcal{L}_{\rm R}$: the second configuration removes the reconstruction loss and trains the model using only the soft contrastive loss.
+Experimental results show that this configuration outperforms the masked modeling approach, indicating that contrastive learning is more effective for EEG-based emotion recognition by capturing high-level semantic features of EEG signals.
+(3) SCMM: the last configuration trains the model with the soft contrastive loss and reconstruction loss, representing the full SCMM model.
+The results demonstrate that this configuration achieves the best performance in all experiments, indicating that SCMM significantly enhances the model performance and stability by combining soft contrastive learning and aggregate reconstruction.
+This improvement is evident under different cross-corpus conditions, demonstrating the feasibility of extending SCMM to real-life aBCI applications.
+Figure 5. Model performance with limited labeled data for fine-tuning on SEED and SEED-IV under same-class and different-class cross-corpus conditions. Best viewed in color. Zoom in for a better view.
+[…]
+Model performance with limited labeled data for fine-tuning on SEED and SEED-IV.
+6.2. Model Performance with Limited Data
+We investigate the model performance of SCMM on the SEED and SEED-IV datasets when fine-tuning with limited labeled data.
+Specifically, we randomly select 1%, 5%, 10%, and 20% of labeled samples from the fine-tuning dataset for model fine-tuning, while the remaining samples are used for testing.
 Figure 5 shows the classification accuracy curves.
@@
 Meanwhile, our model achieves classification accuracies over 60% and 70% with only 1% of labeled data in the SEED3 $\rightarrow$ SEED-IV3 and SEED3 $\rightarrow$ SEED-IV4 experiments.
-The classification accuracies exceed 90% when fine-tuning with 10% of labeled samples in both experiments.
-In summary, experimental results indicate that SCMM maintains superior performance even with limited labeled data for fine-tuning, showing its outstanding robustness and potential in few-shot scenarios.
-The complete results and computational complexity analysis are provided in Appendix F.
-Comparison of different masking strategies
-Strategies SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
-Random 90.30 / 08.80 84.63 / 10.99
-Channel 90.25 / 08.68 85.91 / 10.97
-Hybrid 91.61 / 07.56 87.24 / 08.35
-Table 5: Comparative experiments of different masking strategies on the SEED and SEED-IV 3-category datasets.
-To assess the impact of various masking strategies, we conduct comparative experiments on the SEED and SEED-IV 3-category datasets using three strategies: random, channel, and hybrid masking.
-Table 5 presents the experimental results, showing that the hybrid masking strategy achieves the highest accuracy and the lowest standard deviation in both experiments.
-This suggests that the integration of different masking approaches significantly improves the model performance and stability.
-More details on the different masking strategies are presented in Appendix D.
-Comparison of soft CL in the original data space and embedding space
-While soft contrastive learning has been explored across various domains, most methods focus on computing soft assignments for contrastive loss in the embedding space (Dwibedi et al.
-2021; Yèche et al.
-2021).
-However, we argue that utilizing similarities in the original data space offers superior self-supervision and is particularly well-suited for EEG emotional data.
-To validate this, we conduct additional experiments on the SEED, SEED-IV and DEAP datasets to verify the effectiveness of soft contrastive learning in the original data space.
+The accuracies exceed 90% when fine-tuning with 10% of labeled samples in both experiments.
+In summary, the results indicate that SCMM maintains superior performance even with limited labeled data for fine-tuning, showing its outstanding robustness and potential in few-shot scenarios.
+Figure 6.
+Hyperparameter sensitivity analysis.
+(a) - (f) represent the masking ratio $r$, threshold $\mu$, metric function $Dist(\cdot,\cdot)$, upper bound $\alpha$, sharpness $\tau_{\rm s}$, and temperature $\tau_{\rm c}$, respectively.
+Best viewed in color.
+[…]
+Sensitivity analysis of hyperparameters.
+6.3. Delve into Contrastive Learning
+6.3.1. Hard CL vs. Soft CL
+We compare the model performance of SCMM using traditional hard CL and soft CL on SEED, SEED-IV, and DEAP under same-class and different-class cross-corpus conditions to verify the effectiveness of the proposed soft weighting mechanism.
+Table 4 presents the comparison results, showing that the SCMM model using the soft weighting mechanism consistently achieves the highest classification accuracies and lowest standard deviations in all experiments.
+This improvement indicates that assigning soft weights (rather than hard values) to different sample pairs when computing the contrastive loss effectively captures the inter-sample similarity relationships, thereby improving the discriminative ability of the model.
+6.3.2. Soft CL in the Embedding Space vs. Original Data Space
+While soft CL has been explored across various domains, most methods focus on computing soft assignments for contrastive loss in the embedding space (Dwibedi et al., 2021; Yèche et al., 2021).
+However, we argue that utilizing similarities in the original data space can provide better self-supervision and is particularly suitable for emotional EEG data.
+To validate this, we conduct experiments on SEED, SEED-IV, and DEAP under same-class and different-class cross-corpus conditions to compare soft CL in different spaces.
 Specifically, we modify the metric function $Dist(\cdot,\cdot)$ to use similarities between projected embeddings, shifting the computation of soft assignments from the original data space to the embedding space.
-Table 4 presents the experimental results, demonstrating that soft contrastive learning in the original data space consistently outperforms the embedding space in all experiments.
-Furthermore, this approach allows for pre-computing cosine similarities between original sample pairs offline, thus reducing computational resource requirements and improving training efficiency.
-Visualization
-Figure 6: (a) Heat maps of pairwise similarity matrices. (b) t-SNE visualization of the learned embeddings. Best viewed in color. Zoom in for better view.
-To evaluate whether sample-wise relationships are preserved in the encoder, we randomly select 100 test samples from the SEED dataset, and visualize the pairwise cosine similarity between sample pairs.
-Additionally, we select all test samples of one subject from the SEED dataset and visualize the learned embeddings of SCMM using t-SNE (Van der Maaten and Hinton 2008).
+Table 5 presents the experimental results, demonstrating that soft CL in the original data space consistently outperforms the embedding space in all experiments.
+Furthermore, this approach allows offline pre-computation of cosine similarities of different sample pairs in the original data space, thus reducing computational costs and improving training efficiency.
+6.4. Hyperparameter Sensitivity Analysis
+To analyze the hyperparameter sensitivity of SCMM, we conduct experiments on SEED and SEED-IV under the same-class condition, as shown in Fig.
+6.
+Specifically, the examined hyperparameters are divided into two groups: (1) Hybrid Masking: masking ratio $r$ and threshold $\mu$, and (2) Soft Contrastive Learning: metric function $Dist(\cdot,\cdot)$, upper bound $\alpha$, sharpness $\tau_{\rm s}$, and temperature $\tau_{\rm c}$.
+(1) Hybrid Masking.
+Figures 6(a) - (b) show the classification curves of SCMM with different masking ratios $r$ and thresholds $\mu$.
+Experimental results show that the proposed SCMM achieves the best performance when the masking ratio and threshold are set to $r=0.5$ and $\mu=0.1$ (i.e., the ratio of random masking and channel masking is 9:1), respectively.
+In addition, we find that our model performs better when using hybrid masking compared to random masking or channel masking in most settings, demonstrating the effectiveness of the proposed hybrid masking strategy in enhancing the model performance.
+(2) Soft Contrastive Learning.
+Figures 6(c) - (f) depict the classification curves of SCMM using different metric function $Dist(\cdot,\cdot)$, upper bound $\alpha$, sharpness $\tau_{\rm s}$, and temperature $\tau_{\rm c}$, respectively.
+Specifically, our model performs best on the SEED and SEED-IV 3-category datasets when using cosine similarity as the metric function $Dist(\cdot,\cdot)$, as shown in Fig.
+6(c).
+Additionally, Figures 6(d) - (f) show that SCMM achieves the best performance when the upper bound, sharpness and temperature are set to $\mu=0.5$, $\tau_{\rm s}=0.05$ and $\tau_{\rm c}=0.5$, respectively.
+In summary, the results demonstrate that SCMM maintains excellent performance under different hyperparameter settings, indicating that its superior generalization ability is not significantly affected by hyperparameter changes.
+6.5. Visualization
+6.5.1. Sample-Wise Relationships
+To evaluate whether sample-wise relationships are preserved in the encoder, we randomly select 100 test samples from the SEED dataset and visualize the pairwise cosine similarity between sample pairs.
+Additionally, we select all test samples of one subject from the SEED dataset and visualize the learned embeddings of SCMM using t-SNE (Van der Maaten and Hinton, 2008).
 Figure 7(a) presents heat maps of pairwise similarity matrices, where darker colors indicate higher similarity between samples.
@@
 Moreover, the results of t-SNE visualization in Fig.
-6(b) indicate that the proposed SCMM better clusters samples within the same category and increases the inter-class distance compared to traditional hard CL, thus enhancing the classification performance.
-More visualization results can be found in Appendix G.
-Conclusion
-This paper proposes a novel self-supervised pre-training framework, Soft Contrastive Masked Modeling (SCMM), for cross-corpus EEG-based emotion recognition.
-Unlike traditional contrastive learning models, SCMM integrates soft contrastive learning with a hybrid masking strategy to effectively capture the ”short-term continuity” characteristics inherent in human emotions, and produce stable and generalizable EEG representations.
-Additionally, an aggregator is developed to weightedly aggregate complementary information from multiple close samples, thereby enhancing fine-grained feature representation capability in the modeling process.
-Extensive experiments on three well-recognized datasets show that SCMM consistently achieves SOTA performance in cross-corpus EEG-based emotion recognition tasks under both same-class and different-class conditions.
-Comprehensive ablation study and parameter analysis confirm the superior performance and robustness of SCMM.
-Visualization results indicate that our model effectively reduces the distance between similar samples within the same category, and captures more fine-grained relationships across samples.
+7(b) indicate that our model better clusters samples within the same category and increases the inter-class distance compared to hard CL, thus enhancing the classification performance.
+Figure 7. (a) Heat maps of pairwise similarity matrices. (b) t-SNE visualization of the learned embeddings. Best viewed in color. Zoom in for a better view.
+[…]
+Visualization.
+6.5.2. Intra- and Inter-Class Similarities
+Figure 8.
+Intra- and inter-class cosine similarities of embeddings learned by hard CL and soft CL.
+(a) and (b) represent the average intra-class and inter-class cosine similarities, respectively.
+Best viewed in color.
+[…]
+Intra- and inter-class cosine similarities.
+Figure 9.
+Comparison of reconstruction quality.
+We visualize the reconstruction results to compare the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM).
+Best viewed in color.
+[…]
+Comparison of reconstruction quality.
+To assess the quality of the embeddings learned by SCMM, we randomly select one subject from the SEED dataset and calculate both the average intra- and inter-class cosine similarities between the learned embeddings of all test samples, as shown in Fig.
+8.
+It is evident that the proposed SCMM produces embeddings with higher intra-class similarity compared to traditional hard CL.
+In addition, the average inter-class similarity of the embeddings learned by SCMM is significantly lower than that of hard CL.
+In summary, visualization results confirm that the soft contrastive learning strategy designed in SCMM effectively clusters samples within the same category while distinctly separating samples from different categories, thus enhancing the model’s discriminative capabilities.
+6.5.3. Reconstruction Quality
+To verify the effectiveness of the similarity-aware aggregator designed in SCMM, we compare the reconstruction quality of the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM) on the DEAP dataset.
+For clarity, we flatten both the original input sample and the reconstructed sample into one-dimensional vectors with dimensions $C\times F$ (channels $\times$ features).
+The results depicted in Fig.
+9 illustrate that our model achieves lower reconstruction loss (MSE) and better sample reconstruction, which is crucial to improving the model performance.
+7. Conclusions
+This paper proposes a novel SCMM framework to tackle the critical challenge of cross-corpus generalization in the field of EEG-based emotion recognition.
+Unlike traditional CL-based models, SCMM integrates soft contrastive learning with a hybrid masking strategy to effectively capture the ”short-term continuity” characteristics inherent in human emotions and produce stable and generalizable EEG representations.
+Additionally, a similarity-aware aggregator is introduced to fuse complementary information from semantically related samples, thereby enhancing the fine-grained feature representation capability of the model.
+Extensive experimental results on three well-recognized datasets show that SCMM consistently achieves SOTA performance in cross-corpus EEG-based emotion recognition tasks under both same-class and different-class conditions.
+Comprehensive ablation study and hyperparameter sensitivity analysis confirm the superiority and robustness of SCMM.
+Visualization results indicate that our model effectively reduces the distance between similar samples within the same category and captures more fine-grained relationships across samples.
 These findings suggest that SCMM enhances the feasibility of extending the proposed method to real-life aBCI applications.
+8. Acknowledgements
+This work was funded by the National Natural Science Foundation of China (62276169), the STI 2030-Major Projects (2021ZD0200500), the Medical-Engineering Interdisciplinary Research Foundation of Shenzhen University (2023YG004), the Shenzhen-Hong Kong Institute of Brain Science-Shenzhen Fundamental Research Institutions (2023SHIBS0003), the Open Research Fund of the State Key Laboratory of Brain-Machine Intelligence, Zhejiang University (Grant No.
+BMI2400008), and the Shenzhen Science and Technology Program (No.
+JCYJ20241202124222027 and JCYJ20241202124209011).
 Appendix A Datasets
-A.1 Dataset Description
-We conduct extensive experiments on three well-recognized datasets, SEED (Zheng and Lu 2015), SEED-IV (Zheng et al.
-2018), and DEAP (Koelstra et al.
-2011), to evaluate the model performance of SCMM in cross-corpus EEG-based emotion recognition tasks.
+A.1. Dataset Description
+We conduct extensive experiments on three well-known datasets, SEED (Zheng and Lu, 2015), SEED-IV (Zheng et al., 2018), and DEAP (Koelstra et al., 2011), to evaluate the model performance of SCMM in cross-corpus EEG-based emotion recognition tasks.
+These datasets encompass different EEG equipment, emotional stimuli, data specifications, labeling approaches, and subjects, making them exceptionally suitable for evaluating the model’s efficacy in cross-corpus scenarios.
 Table 6 provides a detailed description of the three datasets.
+Table 6. Detailed descriptions of the experimental datasets.
 Datasets Subjects Sessions $\times$ Trials Channels Sampling Rate Classes
-SEED 15 3 $\times$ 15 62 1kHz 3 (Negative, Neutral, Positive)
-SEED-IV 15 3 $\times$ 15 62 1kHz 4 (Sad, Neutral, Fear, Happy)
-DEAP 32 1 $\times$ 40 32 512Hz Valence: 1 - 9, Arousal: 1 - 9
-Table 6: Detailed description of the experimental datasets.
-(1) SEED (Zheng and Lu 2015) was developed by Shanghai Jiao Tong University.
+SEED (Zheng and Lu, 2015) 15 3 $\times$ 15 62 1 kHz 3 (Negative, Neutral, Positive)
+SEED-IV (Zheng et al., 2018) 15 3 $\times$ 24 62 1 kHz 4 (Sad, Neutral, Fear, Happy)
+DEAP (Koelstra et al., 2011) 32 1 $\times$ 40 32 512 Hz Valence: 1 - 9, Arousal: 1 - 9
+(1) SEED (Zheng and Lu, 2015) was developed by the Center for Brain-like Computing and Machine Intelligence (BCMI) of Shanghai Jiao Tong University.
 The dataset used a 62-channel ESI NeuroScan System based on the international 10-20 system to record EEG signals from 15 subjects (7 males and 8 females) under different video stimuli at a sampling rate of 1 kHz.
 Each subject participated in 3 sessions.
-In each session, each subject was required to watch 15 movie clips consisting of 3 different emotional states: negative, neutral and positive.
+In each session, each subject was required to watch 15 movie clips consisting of 3 different emotional states: negative, neutral, and positive.
 Each emotional state contains a total of 5 movie clips, corresponding to 5 trials.
-(2) SEED-IV (Zheng et al.
-2018) used the same EEG acquisition equipment as the SEED dataset, but with different video stimuli, emotion categories and subjects.
+(2) SEED-IV (Zheng et al., 2018) used the same EEG equipment as the SEED dataset, but with different video stimuli, emotion categories, and subjects.
 The dataset recorded EEG signals from 15 subjects under different video stimuli at a sampling rate of 1 kHz.
 Each subject participated in 3 sessions.
-In each session, each subject was required to watch 24 movie clips containing 4 different emotions: sad, neutral, fear and happy.
+In each session, each subject was required to watch 24 movie clips containing 4 different emotions: sad, neutral, fear, and happy.
 Each emotion contains a total of 6 movie clips, corresponding to 6 trials.
-(3) DEAP (Koelstra et al.
-2011) was constructed by Queen Mary University of London.
-The dataset used a 128-channel Biosemi ActiveTwo System to record EEG signals from specific 32 channels of 32 subjects (16 males and 16 females) while watching 40 one-minute music videos at a sampling rate of 512Hz.
-The 40 one-minute videos elicited different emotions according to the valence-arousal dimension.
-Specifically, the valence-arousal emotional model, first proposed by Russell (Russell 1980), places each emotional state on a two-dimensional scale.
+(3) DEAP (Koelstra et al., 2011) was constructed by Queen Mary University of London.
+The dataset has completely different acquisition device, emotional stimuli, data specifications, labeling approaches, experimental protocols, and subjects from the SEED-series datasets.
+Specifically, the DEAP dataset used a 128-channel Biosemi ActiveTwo System to record EEG signals from specific 32 channels of 32 subjects (16 males and 16 females) while watching 40 one-minute music videos at a sampling rate of 512 Hz.
+The 40 videos elicited different emotions according to the valence-arousal dimension.
+Specifically, the valence-arousal emotional model proposed by Russell (Russell, 1980) places each emotional state on a two-dimensional scale.
 The first dimension represents valence, ranging from negative to positive, and the second dimension represents arousal, ranging from calm to exciting.
 Participants rated valence and arousal using a continuous scale of 1 to 9 after watching each video clip.
-A.2 Pre-processing Procedures
-For the SEED and SEED-IV datasets, the raw EEG signals were initially downsampled to 200Hz and filtered through a bandpass filter of 0.3-50Hz to filter noise and remove artifacts.
-Then, the data were divided into multiple non-overlapping segments using sliding windows of 1 second for SEED and 4 seconds for SEED-IV, respectively.
+A.2. Pre-processing Procedures
+For the SEED and SEED-IV datasets, the raw EEG signals were first downsampled to 200 Hz and filtered through a bandpass filter of 0.3-50 Hz to remove noise and artifacts.
+Then, the data were divided into multiple non-overlapping segments using a sliding window of 1s (SEED) and 4s (SEED-IV), respectively.
 After that, we extracted differential entropy (DE) features for each channel of each segment at five frequency bands: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz).
-Finally, the DE features from 62 channels and 5 bands were formed into a feature matrix of shape $62\times 5$, which serves as input to the SCMM model.
+Finally, the DE features from 62 channels and 5 bands were formed into a feature matrix of shape 62×562$\times$ 562 × 5, which serves as input to the SCMM model.
 The extraction of DE features can be expressed as:
-$DE(x)=\frac{1}{2}{\rm log}(2\pi e\sigma^{2}),$ (9)
-Here, $x$ represents an EEG signal segment of a specific length that approximately obeys a Gaussian distribution $N(\mu,\sigma^{2})$, where $\sigma$ denotes the standard deviation of $x$, and $e$ is the Euler constant.
-For the DEAP dataset, the raw EEG signals were first downsampled to 128Hz and denoised by a bandpass filter of 4-45Hz.
+(9) $DE(x)=\frac{1}{2}{\rm log}(2\pi e\sigma^{2}),$
+where $e$ is the Euler constant.
+$x$ is an EEG signal of a specific length that approximately obeys a Gaussian distribution $N(\mu,\sigma^{2})$.
+Here, $\mu$ and $\sigma$ are the mean and standard deviation of $x$, respectively.
+For the DEAP dataset, the raw EEG signals were initially downsampled to 128 Hz and denoised by a bandpass filter of 4-45 Hz.
 Subsequently, the data were segmented into multiple non-overlapping segments using a sliding window of 1s.
 Similar to the SEED and SEED-IV datasets, DE features were extracted for each channel of each segment at five frequency bands.
-Finally, the DE features from 32 channels and 5 bands were formed into a feature matrix of shape $32\times 5$ as input to the model.
-During the experiments, we divided the continuous labels using a fixed threshold of 5 to convert them to binary classification tasks (low / high).
-A.3 Handling Different Number of Channels
+Finally, the DE features from 32 channels and 5 bands were formed into a feature matrix of shape 32×532$\times$ 532 × 5 as input to the model.
+In the experiments, we divided the continuous labels using a fixed threshold of 5 to convert them to binary classification tasks (low/high).
+A.3. Handling Different Number of Channels
 Since the SEED-series datasets and the DEAP dataset contain different numbers of electrodes (channels), we require channel processing before inputting DE features into the model.
-Specifically, we consider the fine-tuning dataset as an anchor.
+Specifically, we consider the fine-tuning dataset as the standard.
 When the number of channels in the fine-tuning dataset is less than in the pre-training dataset, we select data from the corresponding channels in the pre-training dataset and drop the data from the redundant channels as inputs (e.g., pre-training on SEED and fine-tuning on DEAP).
 Conversely, when the number of channels in the fine-tuning dataset is greater than in the pre-training dataset, we fill the missing channel data with zeros in the pre-training dataset to match the fine-tuning dataset (e.g., pre-training on DEAP and fine-tuning on SEED).
-Appendix B Implementation Details
-To reduce computational load while maintaining model performance, we adopt a lightweight design for each module of SCMM.
-Specifically, we use a 3-layer 1D CNN for the encoder $E$ and a 2-layer MLP for the projector $P$.
-For the lightweight decoder $D$, we utilize a single-layer MLP for reconstruction.
-Regarding hyperparameter selection in the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking, i.e., the percentage of random masking and channel masking is 0.9 and 0.1, respectively.
-We use the negative of cosine similarity as the metric function $Dist(\cdot,\cdot)$, and we set $\alpha$ to 0.5, $\tau_{w}$ to 0.05 and $\tau_{c}$ to 0.5 for soft contrastive learning.
-We use Adam as the optimizer with a learning rate of $5\times 10^{-4}$ and an L2-norm penalty coefficient $3\times 10^{-4}$.
-The pre-training process is conducted over 200 epochs with a batch size of 256.
-We save the model parameters $\theta$ from the final epoch as the pre-trained SCMM.
-In the fine-tuning stage, we input the encoded embeddings $h_{i}$ into a classifier consisting of a 2-layer fully connected network for final emotion recognition.
-The Adam optimizer is utilized with a learning rate of $5\times 10^{-4}$ and a weight decay of $3\times 10^{-4}$.
-The number of fine-tuning epochs is set to 50 for the SEED and SEED-IV datasets and 500 for the DEAP dataset, with a batch size of 128.
-For efficient deployment and testing of the model, the pre-trained SCMM is optimized solely using cross-entropy loss during fine-tuning.
-All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce 3090 GPU.
-   Pre-training    Fine-tuning
+Table 7. Hyperparameter settings of SCMM.
+Hyperparameters Pre-training Fine-tuning
 Encoder 3-layer 1D CNN
@@
 Upper Bound $\alpha$ 0.5 -
-Sharpness $\tau_{w}$ 0.05 -
-Temperature $\tau_{c}$ 0.5 -
+Sharpness $\tau_{\rm s}$ 0.05 -
+Temperature $\tau_{\rm c}$ 0.5 -
 Epoch 200 50, 500
@@
 Batch Size 256 128
-Table 7: Hyperparameter settings of SCMM.
-Appendix C Baseline Methods and Experimental Settings
-C.1 Baseline Methods
-We compare the proposed SCMM against eight competitive methods, including four traditional deep learning methods: BiDANN (Li et al.
-2018), TANN (Li et al.
-2021), PR-PL (Zhou et al.
-2023) and E2STN (Zhou et al.
-2024), as well as four self-supervised learning models: SimCLR (Chen et al.
-2020; Tang et al.
-2021), Mixup (Zhang et al.
-2018; Wickstrøm et al.
-2022), MAE (He et al.
-2022) and JCFA (Liu et al.
-2024a).
-Details of the eight baseline methods are summarized as follows.
-  • BiDANN (Li et al.
-2018): The bi-hemispheres domain adversarial neural network mapped the EEG data of both left and right hemispheres into discriminative feature spaces separately to address domain shifts in EEG-based emotion recognition tasks.
-  • TANN (Li et al.
-2021): The transferable attention neutral network is a novel transfer learning methods which learned the discriminative information from EEG signals using local and global attention mechanisms.
-  • PR-PL (Zhou et al.
-2023): The prototypical representation based pairwise learning framework adopted pairwise learning to model the relative relationships between EEG sample pairs in terms of prototypical representations, addressing the critical issues of individual differences and noise labels in cross-subject scenarios.
-  • E2STN (Zhou et al.
-2024): The emotional EEG style transfer network integrated content information from the source domain with style information from the target domain, achieving superior performance in cross-corpus EEG-based emotion recognition tasks.
-  • SimCLR (Chen et al. 2020; Tang et al. 2021): A seminal work in self-supervised contrastive learning, first proposed for computer vision, has been extended to human activity recognition.
-  • Mixup (Zhang et al.
-2018; Wickstrøm et al.
-2022): A novel CL-based data augmentation method that aimed to correctly predict the mixing proportion of two samples, has been applied to time series analysis.
-  • MAE (He et al.
-2022): A groundbreaking work in the field of mask modeling, which proposed to reconstruct the masked portion based on the unmasked portion of the masked sample, has achieved remarkable success in a wide range of fields.
-  • JCFA (Liu et al.
-2024a): The joint contrastive learning framework performed joint contrastive learning across two domains to synchronize the time- and frequency-based embeddings of the same EEG sample in the latent time-frequency space, achieving state-of-the-art performance in cross-corpus scenarios.
-To ensure a fair comparison, we adopt the same encoder, projector, decoder, and classifier structures for SimCLR, Mixup and MAE as used in SCMM.
-We use the default hyperparameters reported in the original paper for all methods in our experiments, unless otherwise specified.
-Additionally, for BiDANN, TANN, PR-PL, E2STN, MAE, and SCMM, the input samples are preprocessed 1-s DE features.
-In contrast, SimCLR, Mixup, and JCFA use preprocessed 1-s EEG signals as inputs, in accordance with the specific design of each model.
-C.2 Experimental Settings
-We adopt a cross-corpus subject-independent experimental protocol in the experiments, following the setup used in JCFA.
-Specifically, samples from one dataset are used for pre-training, while samples of each subject from another dataset are used individually for fine-tuning and testing.
-During the fine-tuning process, we employ a leave-trials-out setting, where samples from a part of trials in each session of each subject in the fine-tuning dataset are used for fine-tuning, and the remaining trials are used for testing.
-For example, SCMM is pre-trained on the SEED-IV dataset and fine-tuned on the SEED dataset.
-The allocation of samples for fine-tuning/testing is as follows: 9/6 for the SEED dataset, 12/6 for the SEED-IV 3-category dataset, 16/8 for the SEED-IV 4-category dataset, and 24/16 for the DEAP dataset.
-This approach effectively avoids information leakage.
-The detailed experimental settings and data division for pre-training and fine-tuning are summarized in Table 8.
+Table 8. Experimental scenarios and data division for cross-corpus EEG-based emotion recognition.
 Evaluations Scenarios Pre-training Fine-tuning/Testing
-Same-Class SEED-IV3 $\rightarrow$ SEED3 SEED-IV, 3-category SEED: 9 / 6 trials in each session per subject
-SEED3 $\rightarrow$ SEED-IV3 SEED SEED-IV, 3-category: 12 / 6 trials in each session per subject
-Different-Class SEED-IV4 $\rightarrow$ SEED3 SEED-IV, 4-category SEED: 9 / 6 trials in each session per subject
-SEED3 $\rightarrow$ SEED-IV4 SEED SEED-IV, 4-category: 16 / 8 trials in each session per subject
-DEAP $\rightarrow$ SEED3 DEAP SEED: 9 / 6 trials in each session per subject
-SEED3 $\rightarrow$ DEAP (Valence) SEED DEAP (Valence): 24 / 16 trials per subject
-SEED3 $\rightarrow$ DEAP (Arousal) SEED DEAP (Arousal): 24 / 16 trials per subject
-Table 8: Pre-training and fine-tuning scenarios under two conditions for cross-corpus EEG-based emotion recognition.
+Same-Class SEED-IV3 $\rightarrow$ SEED3 SEED-IV, 3-class SEED: 9/6 trials in each session of each subject
+SEED3 $\rightarrow$ SEED-IV3 SEED SEED-IV, 3-class: 12/6 trials in each session of each subject
+Different-Class SEED-IV4 $\rightarrow$ SEED3 SEED-IV, 4-class SEED: 9/6 trials in each session of each subject
+SEED3 $\rightarrow$ SEED-IV4 SEED SEED-IV, 4-class: 16/8 trials in each session of each subject
+DEAP $\rightarrow$ SEED3 DEAP SEED: 9/6 trials in each session of each subject
+SEED3 $\rightarrow$ DEAP (Valence) SEED DEAP (Valence): 24/16 trials of each subject
+SEED3 $\rightarrow$ DEAP (Arousal) SEED DEAP (Arousal): 24/16 trials of each subject
+Appendix B Implementation Details
+To reduce computational load while maintaining model performance, we adopt a lightweight design for each module of SCMM.
+Specifically, we use a 3-layer 1D CNN for the encoder $E$ and a 2-layer MLP for the projector $P$.
+For the lightweight decoder $D$, we utilize a single-layer MLP for reconstruction.
+For the hyperparameter selection in the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking, i.e., the ratio of random masking and channel masking is 9:1.
+We use the negative of cosine similarity as the metric function $Dist(\cdot,\cdot)$, and we set $\alpha$ to 0.5, $\tau_{w}$ to 0.05 and $\tau_{c}$ to 0.5 for soft contrastive learning.
+We use Adam optimizer with an initial learning rate of $5\times 10^{-4}$ and an L2-norm penalty coefficient $3\times 10^{-4}$.
+The pre-training process is conducted over 200 epochs with a batch size of 256.
+We save the model parameters $\theta$ from the final epoch as the pre-trained SCMM.
+In the fine-tuning stage, we input the encoded embeddings $h_{i}$ into an emotion classifier consisting of a 2-layer MLP for final emotion recognition.
+The Adam optimizer is utilized with an initial learning rate of $5\times 10^{-4}$ and a weight decay of $3\times 10^{-4}$.
+The number of fine-tuning epochs is set to 50 for SEED and SEED-IV and 500 for DEAP, with a batch size of 128.
+For efficient deployment and testing of the model, the pre-trained SCMM is optimized solely using cross-entropy loss during fine-tuning.
+All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce 3090 GPU.
+Table 7 summarizes the hyperparameter settings of SCMM.
+Appendix C Baselines and Experimental Setup
+We compare the proposed SCMM against ten competitive baselines, including five transfer learning methods: DANN (Ganin et al., 2016), BiDANN (Li et al., 2018b), TANN (Li et al., 2021), PR-PL (Zhou et al., 2023), and E2STN (Zhou et al., 2025), as well as five self-supervised learning models: SimCLR (Chen et al., 2020; Tang et al., 2020), Mixup (Zhang, 2017; Wickstrøm et al., 2022), TS-TCC (Eldele et al., 2021), MAE (He et al., 2022a), and JCFA (Liu et al., 2024).
+Note that E2STN and JCFA are two state-of-the-art (SOTA) methods designed for cross-corpus EEG-based emotion recognition.
+To ensure a fair comparison, we adopt the same encoder, projector, decoder, and classifier structures as SCMM for SimCLR, Mixup, TS-TCC, and MAE.
+We use the default hyperparameters reported in the original papers for all models in the experiments, unless otherwise specified.
+Additionally, for DANN, BiDANN, TANN, PR-PL, E2STN, MAE, and SCMM, the inputs are preprocessed 1-s DE features.
+In contrast, SimCLR, Mixup, TS-TCC, and JCFA use preprocessed 1-s EEG signals as inputs, in accordance with the specific design of each model.
+[…]
+Figure 10.
+Examples of generated masked samples using different masking strategies.
+The masking ratio is set to $r=0.5$, and the thresholds for hybrid masking are set to $\mu=0.1,0.5$, and 0.90.90.9, respectively.
+Best viewed in color.
+[…]
+Examples of generated masked samples using different masking strategies.
+Table 9. Comparison of different masking strategies on SEED and SEED-IV under the same-class condition.
+Strategies SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
+Random 90.30 / 08.80 84.63 / 10.99
+Channel 90.25 / 08.68 85.91 / 10.97
+Parallel 90.34 / 08.20 86.23 / 09.94
+Hybrid 91.61 / 07.56 87.24 / 08.35
 Appendix D Masking Strategy
@@
 To compare our approach with traditional masking strategies, we explore three different masking rules: random masking, channel masking, and hybrid masking.
-Figure 3 illustrates the difference between these three masking strategies.
-(1) Random Masking: Generate masks using a binomial distribution to randomly mask samples in the feature dimension, setting the values of the masked features to zero.
-(2) Channel Masking: Generate masks using a binomial distribution to randomly mask samples in the channel dimension, setting the values of all features within the masked channels to zero.
-(3) Hybrid Masking: Generate a probability matrix using a uniform distribution that proportionally mixes masks generated from random masking and channel masking.
-Our proposed hybrid masking strategy is highly flexible and can be extended to various datasets by integrating multiple masking strategies in different ratios, which is exceptionally suitable for data with rich semantic information.
-This approach effectively generates more diverse masked samples, encouraging the model to comprehensively capture the inherent relationships of the data.
-Appendix E Hyperparameter Analysis
-We conduct comprehensive experiments to verify the hyperparameter sensitivity of SCMM on the SEED and SEED-IV 3-category datasets.
-The hyperparameters examined include the masking ratio $r$, probability threshold $\mu$, metric function $Dist(\cdot,\cdot)$, upper bound $\alpha$, sharpness $\tau_{w}$, and temperature $\tau_{c}$.
-The complete experimental results are presented in the following sections.
-E.1 Masking Ratio
-Table 9 shows the model performance in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments when using different masking ratios (0.1, 0.25, 0.5, and 0.75).
-Experimental results demonstrate that the proposed SCMM consistently achieves the best performance when the masking ratio is set to $r=0.5$.
-Ratio ACC / STD (%)
-SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
+Figure 10 presents examples of generated masked samples using three strategies.
+  • Random Masking: Generate masks using a binomial distribution to randomly mask samples along the feature dimension, setting the values of masked features to 0.
+  • Channel Masking: Generate masks using a binomial distribution to randomly mask samples along the channel dimension, setting the values of all features within the masked channels to 0.
+  • Hybrid Masking: Generate a probability matrix using a uniform distribution that proportionally mixes masks generated by random masking and channel masking.
+To assess the impact of various masking strategies, we conduct comparative experiments on the SEED and SEED-IV 3-category datasets using four strategies: random, channel, parallel, and hybrid masking.
+Specifically, the parallel masking strategy indicates that each sample is augmented randomly using one of the random masking or channel masking, and a threshold $\mu$ is used to control the probability of the two masking strategies being selected.
+Table 9 presents the experimental results, showing that the hybrid masking strategy achieves the highest accuracy and lowest standard deviation in both experiments.
+This suggests that the integration of different masking approaches significantly improves the model performance and stability.
+In addition, parallel masking increases the richness of augmented samples by combining different strategies, which is slightly better than using a single masking approach.
+However, this strategy makes the model training process unstable, resulting in large standard deviations.
+In summary, our proposed hybrid masking strategy is highly flexible and can be extended to various datasets by integrating multiple masking strategies using different ratios, which is exceptionally suitable for data with rich semantic information.
+This strategy effectively generates more diverse masked samples, encouraging the model to comprehensively capture the inherent relationships of the data.
+Table 10. Subject-dependent and subject-independent EEG-based emotion recognition performance on SEED and SEED-IV.
+Methods Subject-Dependent Subject-Independent
+SEED SEED-IV SEED SEED-IV
+SVM (Suykens and Vandewalle, 1999) 83.99 / 09.72 56.61 / 20.02 56.73 / 16.29 37.99 / 12.52
+TCA (Pan et al., 2010) - - 63.38 / 14.88 37.01 / 10.47
+GFK (Gong et al., 2012) - - 71.31 / 14.09 44.04 / 09.31
+M3D (Luo et al., 2025) - - 84.57 / 09.49 60.94 / 08.84
+Deep Learning
+DBN (Zheng and Lu, 2015) 86.08 / 08.34 66.77 / 07.38 - -
+DANN (Ganin et al., 2016) 83.99 / 09.72 56.73 / 16.29 56.61 / 20.02 37.99 / 12.52
+DAN (Li et al., 2018a) - - 83.81 / 08.56 58.87 / 08.13
+DGCNN (Song et al., 2018) 90.40 / 08.49 69.88 / 16.29 79.95 / 09.02 52.82 / 09.23
+BiDANN (Li et al., 2018b) 92.38 / 07.04 70.29 / 12.63 84.14 / 06.87 65.59 / 10.39
+BiHDM (Li et al., 2020) 93.12 / 06.06 74.35 / 14.09 85.40 / 07.53 69.03 / 08.66
+RGNN (Zhong et al., 2020) 94.24 / 05.95 79.37 / 10.54 85.30 / 06.72 73.84 / 08.02
+PR-PL (Zhou et al., 2023) 94.84 / 09.16 83.33 / 10.61 93.06 / 05.12 81.32 / 08.53
+PGCN (Jin et al., 2024) 96.93 / 05.11 82.24 / 14.85 84.59 / 08.68 73.69 / 07.16
+Self-Supervised Learning
+SimCLR (Chen et al., 2020) 81.79 / 11.15 52.47 / 11.57 63.45 / 15.96 50.07 / 11.17
+MoCo (He et al., 2020) 76.58 / 10.72 49.40 / 10.99 58.26 / 15.05 46.19 / 10.04
+SSL-EEG (Xie et al., 2021) 83.32 / 09.20 63.59 / 19.82 67.52 / 12.73 53.62 / 08.47
+GMSS (Li et al., 2022a) 89.18 / 09.74 65.61 / 17.33 76.04 / 11.91 62.13 / 08.33
+SCMM (P-T) 92.19 / 07.68 81.41 / 08.17 83.84 / 06.22 75.28 / 06.37
+SCMM (E-E) 93.02 / 06.67 81.60 / 08.12 84.88 / 05.85 76.09 / 06.77
+Appendix E Within-Dataset Validation
+To evaluate the model performance of SCMM for EEG-based emotion recognition within a single dataset, we conduct additional experiments on SEED and SEED-IV based on two experimental protocols: subject-dependent and subject-independent.
+In the experiments, we compare the proposed SCMM with three different types of models: machine learning, deep learning, and self-supervised learning.
+Additionally, we explore the model performance of SCMM using two different training strategies: pre-training-fine-tuning (P-T) and end-to-end (E-E).
+Table 10 presents the experimental results, indicating that our model achieves competitive results compared with various advanced deep learning models and significantly outperforms traditional machine learning methods and self-supervised learning models.
+In addition, the SCMM model using the end-to-end training strategy is better than the pre-training-fine-tuning strategy due to the introduction of the emotion classifier for joint learning.
+In summary, comprehensive experimental results demonstrate the effectiveness of the proposed SCMM in within-dataset EEG-based emotion recognition tasks.
+Appendix F Full Results
+Tables 11 to 14 present the full experimental results of SCMM.
+Table 11. Full results on SEED, SEED-IV, and DEAP under same-class and different-class cross-corpus conditions.
+Scenarios Accuracy Precision Recall F1 Score AUROC AUPRC
+Same-Class
+SEED-IV3 $\rightarrow$ SEED3 91.61 / 07.56 93.38 / 05.71 91.35 / 07.89 91.02 / 08.40 95.45 / 05.82 91.90 / 09.71
+SEED3 $\rightarrow$ SEED-IV3 87.24 / 08.35 87.10 / 09.02 85.58 / 09.59 84.39 / 10.07 89.91 / 07.78 86.81 / 10.12
+Different-Class
+SEED-IV4 $\rightarrow$ SEED3 91.26 / 07.91 92.62 / 06.79 91.02 / 08.14 90.77 / 08.69 95.59 / 06.12 92.28 / 09.46
+SEED3 $\rightarrow$ SEED-IV4 80.89 / 08.69 78.02 / 08.39 82.84 / 06.68 77.08 / 07.72 89.98 / 06.09 83.96 / 08.68
+DEAP3 $\rightarrow$ SEED3 91.70 / 08.07 92.94 / 06.85 91.48 / 08.29 91.15 / 08.88 95.45 / 06.64 92.34 / 10.49
+SEED3 $\rightarrow$ DEAP (Valence) 73.96 / 06.75 68.11 / 08.89 65.77 / 07.75 65.68 / 08.75 74.15 / 07.34 72.12 / 07.68
+SEED3 $\rightarrow$ DEAP (Arousal) 72.66 / 05.67 70.97 / 05.91 68.90 / 07.11 68.84 / 07.57 75.77 / 07.33 74.79 / 07.50
+Table 12. Full results of ablation study on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
+Scenarios Accuracy Precision Recall F1 Score AUROC AUPRC
+w/o Soft Contrastive Loss $\mathcal{L}_{\rm C}$
+SEED-IV3 $\rightarrow$ SEED3 89.68 / 09.32 91.81 / 07.76 89.39 / 09.58 89.20 / 09.97 94.71 / 05.95 91.73 / 09.32
+SEED3 $\rightarrow$ SEED-IV3 84.24 / 11.90 84.79 / 14.99 83.84 / 10.79 81.94 / 14.11 90.65 / 10.10 85.69 / 15.93
+SEED-IV4 $\rightarrow$ SEED3 89.45 / 09.10 91.23 / 08.34 89.21 / 09.32 89.13 / 09.58 95.11 / 05.86 91.24 / 10.77
+SEED3 $\rightarrow$ SEED-IV4 77.24 / 09.14 73.21 / 14.36 73.55 / 12.89 70.24 / 14.29 88.57 / 09.82 80.79 / 16.07
+w/o Reconstruction Loss $\mathcal{L}_{\rm R}$
+SEED-IV3 $\rightarrow$ SEED3 90.73 / 08.48 93.22 / 06.17 90.46 / 08.76 90.26 / 09.07 94.57 / 06.58 90.52 / 10.06
+SEED3 $\rightarrow$ SEED-IV3 85.07 / 11.05 85.26 / 14.01 81.59 / 13.58 80.88 / 15.46 88.35 / 15.99 85.16 / 18.10
+SEED-IV4 $\rightarrow$ SEED3 90.96 / 08.36 92.43 / 07.24 90.70 / 08.57 90.41 / 09.07 94.36 / 06.39 89.98 / 10.40
+SEED3 $\rightarrow$ SEED-IV4 78.32 / 07.19 73.07 / 14.36 74.59 / 12.07 72.01 / 13.92 87.02 / 12.49 82.09 / 16.17
+Full Model
+SEED-IV3 $\rightarrow$ SEED3 91.61 / 07.56 93.38 / 05.71 91.35 / 07.89 91.02 / 08.40 95.45 / 05.82 91.90 / 09.71
+SEED3 $\rightarrow$ SEED-IV3 87.24 / 08.35 87.10 / 09.02 85.58 / 09.59 84.39 / 10.07 89.91 / 07.78 86.81 / 10.12
+SEED-IV4 $\rightarrow$ SEED3 91.26 / 07.91 92.62 / 06.79 91.02 / 08.14 90.77 / 08.69 95.59 / 06.12 92.28 / 09.46
+SEED3 $\rightarrow$ SEED-IV4 80.89 / 08.69 78.02 / 08.39 82.84 / 06.68 77.08 / 07.72 89.98 / 06.09 83.96 / 08.68
+Table 13. Full results of model performance with limited labeled data for fine-tuning on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
+Scenarios Accuracy Precision Recall F1 Score AUROC AUPRC
+1% of Labeled Data
+SEED-IV3 $\rightarrow$ SEED3 74.98 / 17.00 74.52 / 21.19 74.69 / 17.11 70.64 / 21.53 93.76 / 06.20 89.32 / 09.68
+SEED-IV4 $\rightarrow$ SEED3 73.20 / 16.09 73.41 / 21.28 72.93 / 16.23 69.14 / 20.07 94.32 / 04.69 90.31 / 07.53
+SEED3 $\rightarrow$ SEED-IV3 71.35 / 16.29 72.22 / 18.40 69.37 / 15.47 66.23 / 18.19 80.18 / 14.09 74.07 / 17.64
+SEED3 $\rightarrow$ SEED-IV4 60.59 / 23.72 55.23 / 29.74 58.65 / 22.49 52.48 / 27.81 82.19 / 13.44 70.54 / 19.49
+5% of Labeled Data
+SEED-IV3 $\rightarrow$ SEED3 95.23 / 06.98 96.20 / 04.72 95.10 / 07.14 94.86 / 07.88 99.67 / 00.38 99.45 / 00.62
+SEED-IV4 $\rightarrow$ SEED3 90.51 / 12.43 92.90 / 08.38 90.27 / 12.70 89.04 / 15.29 98.23 / 02.25 96.76 / 04.34
+SEED3 $\rightarrow$ SEED-IV3 87.20 / 16.12 88.31 / 14.88 86.08 / 15.69 85.35 / 17.64 91.93 / 10.70 89.14 / 13.71
+SEED3 $\rightarrow$ SEED-IV4 86.01 / 09.76 84.86 / 14.77 82.69 / 11.35 81.29 / 13.75 95.14 / 03.37 92.56 / 06.89
+10% of Labeled Data
+SEED-IV3 $\rightarrow$ SEED3 99.86 / 00.31 99.87 / 00.30 99.86 / 00.31 99.86 / 00.31 100.00 / 00.00 100.00 / 00.00
+SEED-IV4 $\rightarrow$ SEED3 99.72 / 00.80 99.74 / 00.74 99.71 / 00.81 99.72 / 00.80 99.87 / 00.46 99.78 / 00.77
+SEED3 $\rightarrow$ SEED-IV3 90.04 / 16.94 91.00 / 17.05 90.09 / 15.87 89.32 / 18.21 94.29 / 10.86 92.55 / 13.98
+SEED3 $\rightarrow$ SEED-IV4 94.43 / 05.60 95.66 / 04.71 92.18 / 08.23 92.19 / 08.82 98.12 / 02.13 96.03 / 05.45
+20% of Labeled Data
+SEED-IV3 $\rightarrow$ SEED3 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00
+SEED-IV4 $\rightarrow$ SEED3 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00
+SEED3 $\rightarrow$ SEED-IV3 94.16 / 08.52 95.20 / 06.36 93.66 / 08.00 93.72 / 08.35 97.59 / 03.38 96.53 / 04.85
+SEED3 $\rightarrow$ SEED-IV4 98.91 / 02.62 98.82 / 02.97 98.41 / 03.73 98.52 / 03.65 98.97 / 03.20 98.13 / 05.64
+Table 14. Full results of hyperparameter sensitivity analysis on SEED and SEED-IV under the same-class cross-corpus condition.
+Hyperparameters SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
+Masking Ratio $r$
 0.1 91.19 / 08.07 85.05 / 11.68
@@
 0.75 90.24 / 08.08 83.49 / 13.19
-Table 9: Hyperparameter sensitivity analysis of masking ratio $r$ on SEED and SEED-IV.
-E.2 Probability Threshold
-Table 10 presents the model performance using different probability thresholds (ranging from 0 to 1) for hybrid masking.
-Here, $\mu=0$ refers to the exclusive use of the random masking strategy, while $\mu=1$ indicates the exclusive use of the channel masking strategy.
-Experimental results show that SCMM achieves the best performance in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments when the probability threshold is set to $\mu=0.1$ (i.e., the ratio of random masking and channel masking is 9:1).
-Threshold ACC / STD (%)
-SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
+Threshold $\mu$
 0 (Random) 90.30 / 08.80 84.63 / 10.99
@@
 0.8 89.52 / 08.29 86.10 / 12.48
-0.9 89.42 / 08.18 86.05 / 11.75
+0.9 90.25 / 08.68 85.91 / 10.97
 1 (Channel) 90.25 / 08.68 85.91 / 10.97
-Table 10: Hyperparameter sensitivity analysis of threshold $\mu$ on SEED and SEED-IV.
-E.3 Metric Function
-We evaluate the model performance of SCMM using different metric functions in soft assignments $w(\cdot,\cdot)$.
-Experimental results in Table 11 indicate that SCMM performs best on the SEED and SEED-IV 3-category datasets when cosine similarity is used as the metric function.
-Metrics ACC / STD (%)
-SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
+Metric Function $Dist(\cdot,\cdot)$
 Manhattan 90.09 / 09.06 85.83 / 11.14
@@
 Cosine 91.61 / 07.56 87.24 / 08.35
-Table 11: Hyperparameter sensitivity analysis of metric function $Dist(\cdot,\cdot)$ on SEED and SEED-IV.
-E.4 Upper Bound
-We assess the impact of different upper bounds on calculating soft assignments $w(\cdot,\cdot)$.
-To restrict the soft assignments to the range of 0 to 1, we explore the model performance of SCMM with upper bounds of 0.25, 0.5, 0.75, and 1.
-Experimental results in Table 12 illustrate that the proposed SCMM achieves the best performance in the SEED-IV3 $\rightarrow$ SEED3 experiment when the upper bound is set to $\alpha=0.5$.
-Upper Bound ACC / STD (%)
-SEED-IV3 $\rightarrow$ SEED3
-0.25 91.21 / 08.54
-0.5 91.61 / 07.56
-0.75 91.12 / 08.25
-1 90.50 / 08.00
-Table 12: Hyperparameter sensitivity analysis of upper bound $\alpha$ in the SEED-IV3 $\rightarrow$ SEED3 experiment.
-E.5 Sharpness
-Table 13 explore the impact of different sharpness parameters in the SEED-IV3 $\rightarrow$ SEED3 experiment, as shown in Table 13.
-Experimental results indicate that our model achieves the best performance with a sharpness parameter of $\tau_{w}=0.05$.
-Sharpness ACC / STD (%)
-SEED-IV3 $\rightarrow$ SEED3
-0.01 90.93 / 09.13
-0.05 91.61 / 07.56
-0.1 91.08 / 09.00
-0.2 90.59 / 07.40
-0.5 90.63 / 09.16
-1 90.09 / 08.62
-2 89.58 / 09.26
-Table 13: Hyperparameter sensitivity analysis of sharpness $\tau_{w}$ in the SEED-IV3 $\rightarrow$ SEED3 experiment.
-E.6 Temperature
-We conduct comparison experiments on the SEED and SEED-IV 3-category datasets to explore the impact of temperature parameter on model performance.
-Experimental results in Table 14 show that SCMM consistently achieves the best performance in both experiments when the temperature parameter is set to $\tau_{c}=0.5$.
-Temperature ACC / STD (%)
-SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
+Upper Bound $\alpha$
+0.25 91.21 / 08.54 -
+0.5 91.61 / 07.56 -
+0.75 91.12 / 08.25 -
+1 90.50 / 08.00 -
+Sharpness $\tau_{\rm s}$
+0.01 90.93 / 09.13 -
+0.05 91.61 / 07.56 -
+0.1 91.08 / 09.00 -
+0.2 90.59 / 07.40 -
+0.5 90.63 / 09.16 -
+1 90.09 / 08.62 -
+2 89.58 / 09.26 -
+Temperature $\tau_{\rm c}$
 0.05 90.97 / 08.17 85.95 / 11.14
@@
 2 90.62 / 08.66 86.73 / 10.18
-Table 14: Hyperparameter sensitivity analysis of temperature $\tau_{c}$ on SEED and SEED-IV.
-F Model Performance and Computational Complexity Analysis of SCMM
-Metrics 1% of labeled data
-SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3 SEED-IV4 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV4
-ACC / STD (%) 74.98 / 17.00 71.35 / 16.29 73.20 / 16.09 60.59 / 23.72
-Time Cost (s) 232.35 97.11 232.35 130.47
-5% of labeled data
-ACC / STD (%) 95.23 / 06.98 87.20 / 16.12 90.51 / 12.43 86.01 / 09.76
-Time Cost (s) 240.48 100.23 240.48 132.90
-10% of labeled data
-ACC / STD (%) 99.86 / 00.31 90.04 / 16.94 99.72 / 00.80 94.43 / 05.60
-Time Cost (s) 250.81 102.03 250.81 135.67
-20% of labeled data
-ACC / STD (%) 100.0 / 00.00 94.16 / 08.52 100.0 / 00.00 98.91 / 02.62
-Time Cost (s) 267.21 105.48 267.21 139.31
-Model Parameters: 1.52 M
-Table 15: Model performance and computational complexity of SCMM when fine-tuning with limited labeled data (1%, 5%, 10%, and 20%) on SEED and SEED-IV.
-To investigate the trade-off between model performance and computational complexity in few-shot scenarios, we assess the classification accuracy, time cost, and the number of trainable parameters of SCMM when fine-tuning with limited labeled data.
-Experimental results in Table 15 show that the model performance of SCMM significantly improves as the number of fine-tuning labeled samples increases.
-Meanwhile, the inference time cost remains low due to our lightweight design.
-Additionally, our model achieves superior classification performance in cross-corpus EEG-based emotion recognition tasks with very few parameters.
-In summary, the proposed SCMM effectively balances model performance and computational complexity.
-G Visualization
-G.1 Intra- and Inter-Class Similarities
-Figure 7: Intra- and inter-class cosine similarities of embeddings learned by hard CL and soft CL.
-(a) is the average intra-class cosine similarity, and (b) is the average inter-class cosine similarity.
-Best viewed in color.
-To assess the quality of the embeddings learned by SCMM, we randomly select one subject from the SEED dataset, and calculate both the average intra- and inter-class cosine similarities between the learned embeddings of all test samples, as shown in Fig.
-7.
-It is evident that the proposed SCMM produces embeddings with higher intra-class similarity compared to traditional hard contrastive learning.
-In addition, the average inter-class similarity of the embeddings learned by SCMM is significantly lower than that of hard CL.
-In summary, visualization results confirm that the soft contrastive learning strategy designed in SCMM effectively clusters samples within the same category while distinctly separating samples from different categories, thus enhancing the model’s discriminative capabilities.
-G.2 Reconstruction Quality
-Figure 8: Comparison of reconstruction quality.
-We visualize the reconstruction results to compare the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM).
-This case is from the DEAP dataset.
-Best viewed in color.
-To verify the effectiveness of the aggregator designed in SCMM, we compare the reconstruction quality of the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM) on the DEAP dataset.
-For clarity, we flatten both the original input sample and the reconstructed sample into one-dimensional vectors with dimensions $C\times F$ (channels $\times$ features).
-The results depicted in Fig.
-8 illustrate that our model achieves lower reconstruction loss (MSE) and better sample reconstruction.
+Pre-training Batch Size
+8 91.41 / 08.42 86.64 / 11.77
+16 91.29 / 08.90 86.33 / 12.04
+32 91.18 / 08.23 86.77 / 12.72
+64 90.35 / 08.37 86.85 / 10.96
+128 91.00 / 08.22 85.74 / 13.42
+256 91.61 / 07.56 87.24 / 08.35
+512 90.04 / 08.04 86.74 / 11.14